### v2

# Universal Sequence Preconditioning

Abstract
We study the problem of preconditioning in the setting of sequential prediction. From the theoretical lens of linear dynamical systems, we show that applying a convolution to the input sequence translates to applying a polynomial to the unknown transition matrix in the hidden space. With this insight, we develop a novel preconditioning method that convolves the input sequence with the coefficients of the Chebyshev or Legendre polynomials.
We formally prove that this improves the regret of two distinct prediction methods. Moreover, using this preconditioning technique on either method gives the first sublinear regret bounds that are also hidden dimension independent (up to logarithmic factors) even when the hidden transition matrix is asymmetric.
From rigorous experiments on synthetic data we show that our simple preconditioning method generalizes to both 1) settings where the data is not from a linear dynamical system and 2) a broad range of learning algorithms, including recurrent neural networks.
1 Introduction
In sequence prediction the goal of the learner is to predict the next token accurately according to a predetermined loss function, such as the mean square error, or perplexity / cross-entropy loss. This fundamental problem in machine learning has gained increased importance with the rise of large language models, that predict tokens from a dictionary while trying to minimize the cross entropy loss.
The focus of this paper are not methods for sequence prediction, but rather preconditioning algorithms - methods that modify the input sequence and make it more amenable to accurate learning. As an example, consider the widely-used methodology of differencing, proposed by Box and Jenkis in the 70s box1976time , or converting an online sequence $\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{t},...$, to the sequence of differences / deltas,
$\mathbf{y}_{1}-\mathbf{y}_{0}\ ,\ \mathbf{y}_{2}-\mathbf{y}_{1}\ ,\ ...\ ,\ \mathbf{y}_{t}-\mathbf{y}_{t-1}\ ,\ ...$
It is widely acknowledged that learning this sequence is “easier" than learning the original sequence for a large number of modalities. But in what sense is this difference sequence easier to learn, and can we generalize this pre-processing to incorporate more sophisticated techniques with provable guarantees? The question we ask is
What is the general form of sequence preconditioning that enables provable accurate learning?
The more general form of preconditioning, naturally, is one that takes not only the last observation, but last $n$ observations. What is the optimal way to combine these past combinations to make the sequence easier to learn?
Our answer takes us to the theory of dynamical systems. In this domain, we prove that certain preconditioning techniques can greatly improve upon previous prediction methods giving rise to new results. Through the lens of linear dynamical systems this has a beautiful pattern: the preconditioning filter forms coefficients of an $n$ degree polynomial, and the hidden system transition matrix is evaluated on this polynomial– potentially shrinking the domain This allows us to prove the first dimension-independent sublinear regret bounds for asymmetric linear dynamical systems that are marginally stable.
1.1 Our results
Our main contribution is a novel universal sequence preconditioning (USP) method, formally given in Algorithm 1 when the preconditioning coefficients $\mathbf{c}_{0:n}$ are set to be the coefficients of the $n$-th monic Chebyshev polynomial. We provide a theoretical analysis of this algorithm in the online sequence prediction setting when the original algorithm $\mathcal{A}$ is 1) simply regression and 2) spectral filtering. In Algorithm 1 we provide the offline train-test version of the method and in Algorithm 2 we provide the online version.
Algorithm 1 Universal Sequence Preconditioning (Offline Version)
1:  Training
2:  Input: training data $(\mathbf{u}^{1:N}_{1:T},\mathbf{y}^{1:N}_{1:T})$ where $(\mathbf{u}^{i}_{t},\mathbf{y}^{i}_{t})$ is the $t$-th input/output pair in the $i$-th example sequence; preconditioning coefficients $\mathbf{c}_{0:n}$; algorithm $\mathcal{A}$.
3:  Assert that $\mathbf{c}_{0}=1$.
4:  for $i=1$ to $N$ do
[…]
6:  end for
7:  Train algorithm $\mathcal{A}$ on preconditioned training data $\left(\mathbf{u}^{1:N}_{1:T},\mathbf{y}^{\mathrm{preconditioned},1:N}_{1:T}\right)$.
8:  Test Time
9:  for $t=1$ to $T$ do
10:     Receive $\mathbf{u}_{t}$.
11:     Output prediction of $\mathbf{y}_{t}$ as $\hat{\mathbf{y}}_{t}\leftarrow\mathcal{A}\left(\mathbf{u}_{1:t},\mathbf{y}_{1:(t-1)}\right)-\sum_{i=1}^{n}\mathbf{c}_{i}\mathbf{y}_{t-i}.$
12:     Receive $\mathbf{y}_{t}$.
13:  end for
Algorithm 2 Universal Sequence Preconditioning (Online Version)
1:  Input: sequence prediction model $f_{\theta}$ with parameters $\theta$; loss function $\ell(\cdot)$; preconditioning coefficients $\mathbf{c}_{1:n}$.
2:  for $t=1$ to $T$ do
3:     Receive $\mathbf{u}_{t}$
4:     Predict
$\hat{\mathbf{y}}_{t}=-\sum_{j=1}^{n}\mathbf{c}_{j}\mathbf{y}^{i}_{t-j}+f_{\theta^{t}}(\mathbf{u}_{1:t},\mathbf{y}_{1:t-1})$
5:     Observe true output $\mathbf{y}_{t}$ and suffer loss $\ell_{t}(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})$.
6:     Update via projected gradient descent
$\theta^{t+1}\leftarrow\mathrm{proj}_{\mathcal{K}}\left(\theta^{t}-\eta_{t}\nabla\mkern-2.5mu_{\theta}\ell_{t}(\hat{\mathbf{y}}_{t},\mathbf{y}_{t})\right)$
7:  end for
For this universal preconditioning method we give to applications that give new results in sequence prediction as follows.
  1. 1.
    Assuming the signal is generated by a linear dynamical system, we show that our preconditioning method gives rise to the first hidden-dimension free regret bounds for asymmetric dynamical systems that are marginally stable. We show this for two methods: simple regression (Algorithm 3) and spectral filtering (Algorithm 4). We give the formal performance guarantees in Theorem 2.1 (for Algorithm 3) and Theorem C.2 (for Algorithm 4).
  2. 2.
    We evaluate the universal preconditioning algorithm on a variety of signals and sequence prediction methods. Across the board, for the different signal generation as well as recovery technique, universal preconditioning shows improved sequence prediction as compared to other approaches, including learned preconditioning. Our experimental results are in Section 5.
1.2 Related work
Our manuscript is technically involved and incorporates linear dynamical systems, spectral filtering, complex Chebyshev and Legendre polynomials, Hankel and Toeplitz matrix eigendecay, Gaussian quadrature and other techniques. The related work is thus expansive, and due to space limitations we give a detailed treatment in Appendix A.
Preconditioning in the context of time series analysis has roots in the classical work of Box and Jenkins box1976time . In their foundational text they propose differencing as a method for making the time series stationary, and thus amenable to statistical learning techniques such as ARMA (auto-regressive moving average) anava2013online . The differencing operator can be applied numerous times, and for different lags, giving rise to the ARIMA family of forecasting models. Identifying the order of an ARIMA model, and in particular the types of differencing needed to make a series stationary, is a hard problem This is a special case of the problem we consider: differencing corresponds to certain coefficients of preconditioning the time series, whereas we consider arbitrary coefficients.
For a thorough introduction to modern control theory and exposition on open loop / closed loop predictors, learning via regression, and spectral filtering, see hazan2022introduction .
2 Main Results
In this section we formally state our main algorithms and theorems. We show that the Universal Sequence Preconditioning method provides significantly improved regret bounds for learning linear dynamical systems than previously known when used in conjunction with two distinct methods. The first method is simple convex regression, and the second is spectral filtering.
Both algorithms allow for learning in the case of marginally stable linear dynamical systems and allow for certain asymmetric transition matrices of arbitrary high hidden dimension. The regret bounds are free of the hidden dimension (up to logarithmic factors) – which significantly extends the state of the art. After stating the results formally, we explain their significance in context of known results in Section 3.
2.1 Regression Algorithm
Algorithm 3 is an instantiation of Algorithm 2 for the method of convex regression. We set the preconditioning coefficients to be the coefficients of the $n$-th degree (monic) Chebyshev polynomial.
Algorithm 3 Convolutional Preconditioned Regression
1:  Input: initial parameter $\mathbf{Q}^{0}$; preconditioning coefficients $\mathbf{c}_{0:n}$ from the $n$-th degree (monic) Chebyshev polynomial; convex constraints $\mathcal{K}=\left\{\left(\mathbf{Q}_{0},\dots,\mathbf{Q}_{n-1}\right)\textrm{ s.t. }\|\mathbf{Q}_{j}\|~{}\leq~{}C_{\textrm{domain}}\|\mathbf{c}\|_{1}\right\}$
2:  Assert that $\mathbf{c}_{0}=1$.
3:  for $t=1$ to $T$ do
4:     Receive $\mathbf{u}_{t}$
5:     Predict $\hat{\mathbf{y}}_{t}(\mathbf{Q}^{t})=-\sum_{i=1}^{n}c_{i}\mathbf{y}_{t-i}+\sum_{j=0}^{n}\mathbf{Q}^{t}_{j}\mathbf{u}_{t-j}$.
6:     Observe true output $\mathbf{y}_{t}$ and suffer loss $\ell_{t}(\mathbf{Q}^{t})=\|\hat{\mathbf{y}}_{t}(\mathbf{Q}^{t})-\mathbf{y}_{t}\|_{1}$.
7:     Update and project
$\mathbf{Q}^{t+1}\leftarrow\mathrm{proj}_{\mathcal{K}}\left(\mathbf{Q}^{t}-\eta_{t}\nabla\mkern-2.5mu_{\mathbf{Q}}\ell_{t}(\mathbf{Q}^{t})\right)$
8:  end for
Algorithm 3 comes with a formal performance guarantee as stated in the following theorem. For exposition and notation on linear dynamical systems, which are referenced in the theorem see section A.1.
This theorem states vanishing loss compared to the optimal ground-truth predictor, at a rate that is independent of the hidden dimension of the system or its condition number. For the significance of this result, see Section 3.
Theorem 2.1.
Let $\left\{\mathbf{u}_{t}\right\}_{t=1}^{T}\in{\mathcal{R}}^{d_{\textrm{in}}}$ be any sequence of inputs which satisfy $\|\mathbf{u}_{t}\|_{2}~{}\leq~{}1$ and let $\left\{\mathbf{y}_{t}\right\}_{t=1}^{T}$ be the corresponding output coming from some linear dynamical system $(\mathbf{A},\mathbf{B},\mathbf{C})$ as defined per Eq. 3 in Section A.1. Assume that $\|\mathbf{B}\|\|\mathbf{C}\|~{}\leq~{}C_{\textrm{domain}}$. Let $\lambda_{1},\dots,\lambda_{d_{h}}$ denote the spectrum of $\mathbf{A}$. If
$\max_{j\in[d_{h}]}|\arg(\lambda_{j})|~{}\leq~{}1/(64\log_{2}(T))^{2},$
then the predictions $\hat{\mathbf{y}}_{1},\dots,\hat{\mathbf{y}}_{T}$ from Algorithm 3 where the preconditioning coefficients $\mathbf{c}_{0:n}$ are chosen to be the coefficients of the $n$-th monic Chebyshev polynomial satisfy
$\frac{1}{T}\sum_{t=1}^{T}\|\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\|_{1}~{}\leq~{}\frac{9C_{\textrm{domain}}\sqrt{d_{\textrm{out}}}\log_{2}^{2}(3T)}{T^{2/13}}.$
The proof of Theorem 2.1 is in Appendix B.
2.2 Spectral Filtering
Our second main result is an efficient algorithm called Chebyshev Spectral Filtering, which is an instantiation of Algorithm 2 using the Chebyshev coefficients and incorporating a new variant of the spectral filtering algorithm. The algorithm uses a new spectral filtering basis, for which we define two key quantities. Given horizon $T$ and $\alpha\in\mathbb{C}$ let
$\mu_{T}(\alpha)\stackrel{{\scriptstyle\text{def}}}{{=}}\begin{bmatrix}1&\alpha&\dots&\alpha^{T-1}\end{bmatrix}^{\top},$ (1)
and
[…]
where $\overline{\alpha}\in\mathbb{C}$ denotes the complex conjugate. The novel spectral filters are the eigenvectors of $\mathbf{Z}_{T-n-1}$, which we denote as $\phi_{1},\dots,\phi_{T-n-1}$. Note that in the standard spectral filtering literature, the spectral filtering matrix is an integral over the real line and does not involve the complex conjugate. Our new matrix has an entirely different structure and although it looks quite similar, it surprisingly upends the proof techniques to ensure exponential spectral decay, a critical property for the method. We therefore introduce a novel proof of spectral decay in Theorem C.7 found in Section C.1.1.
Algorithm 4 Convolutional Preconditioned Spectral Filtering
1:  Input: initial $\mathbf{Q}^{1}_{1:n},\mathbf{M}^{1}_{1:k}$, horizon $T$, convex constraints $\mathcal{K}=\left\{(\mathbf{Q}_{0},\dots,\mathbf{Q}_{n-1},\mathbf{M}_{1},\dots,\mathbf{M}_{k}\textrm{ s.t. }\|\mathbf{Q}_{j}\|~{}\leq~{}R_{Q}\textrm{ and }\|\mathbf{M}_{j}\|~{}\leq~{}R_{M}\right\}$, parameter $n$, coefficients $\mathbf{c}_{1:n}$.
2:  Let $p_{n}^{\mathbf{c}}(x)=\mathbf{c}_{0}x^{n}+\mathbf{c}_{1}x^{n-1}+\dots+\mathbf{c}_{n}$.
3:  Let $\phi_{1},...,\phi_{n}$ be the top $n$ eigenvectors of $\mathbf{Z}_{T-n-1}$.
4:  Assert $\mathbf{c}_{0}=1$.
5:  for $t=1$ to $T$ do
6:     Let $\tilde{\mathbf{u}}_{t-n-1:1}$ be $\mathbf{u}_{t-n-1:1}$ padded with zeros so it has dimension $T-n-1\times d_{\textrm{in}}$.
7:     Predict $\hat{\mathbf{y}}_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})=-\sum_{i=1}^{n}\mathbf{c}_{i}\mathbf{y}_{t-i}+\sum_{j=0}^{n}\mathbf{Q}^{t}_{j}\mathbf{u}_{t-j}+\frac{1}{\sqrt{T}}\sum_{j=1}^{k}\mathbf{M}_{j}^{t}\phi_{j}^{\top}\tilde{\mathbf{u}}_{t-n-1:1}$.
8:     Observe true $\mathbf{y}_{t}$, define loss $\ell_{t}(\hat{\mathbf{y}}_{t})=\|\hat{\mathbf{y}}_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})-\mathbf{y}_{t}\|_{1}$.
9:     Update and project: $(\mathbf{Q}^{t+1},\mathbf{M}^{t+1})=\mathrm{proj}_{\mathcal{K}}\left(\mathbf{Q}^{t},\mathbf{M}^{t})-\eta_{t}\nabla\mkern-2.5mu\ell_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})\right)$
10:  end for
Two important parameters in our result are as follows.
  1. 1.
    We denote the largest argument of any eigenvalue of the transition matrix $\mathbf{A}$ by
    $\beta=\arg(\lambda_{\max}(\mathbf{A})).$
    This important parameter measures how far away $\mathbf{A}$ is from being symmetric, and as explained in Section A.2, the memory capacity of the system. For the permutation system with memory $d$, it holds that $\beta$ is as large as $1-\frac{1}{d}$.
  2. 2.
    We denote the number of closed-loop components $\mathbf{Q}_{1:n}$ by $n$ and the maximum magnitude of coefficients $c_{1:n}$ by $r$. The parameter $n$ and the coefficients $c_{1:n}$ bounded by $r$ can be chosen to be independent of the hidden dimension. Instead we show that $n$ can be chosen to depend logarithmically with $T$ and $r$ can depend sublinearly on $T$, giving sublinear regret bounds.
We prove that Algorithm 4, instantiated with the coefficients of the Chebyshev polynomial of degree $n=\log T$, achieves regret
$\mbox{{Regret}}_{T}(\mathrm{CSF})=\tilde{O}\left(T^{10/13}\right)$
whenever $\beta<\frac{1}{64n^{2}}$.
A formal statement of the regret guarantee is given in Section C, Theorem C.2.
3 Significance of Our Results
Linear dynamical systems (LDS) are perhaps the most basic and well studied dynamical systems in engineering and control science. Given input vectors $\mathbf{u}_{1},\dots,\mathbf{u}_{T}\in{\mathcal{R}}^{d_{\textrm{in}}}$, the system generates a sequence of output vectors $\mathbf{y}_{1},\dots\mathbf{y}_{T}\in{\mathcal{R}}^{d_{\textrm{out}}}$ according to the law
$\mathbf{x}_{t+1}$ $=\mathbf{A}\mathbf{x}_{t}+\mathbf{B}\mathbf{u}_{t}$ (3)
$\mathbf{y}_{t}$ $=\mathbf{C}\mathbf{x}_{t}+\mathbf{D}\mathbf{u}_{t},$ (4)
where $\mathbf{x}_{0},\dots,\mathbf{x}_{T}\in{\mathcal{R}}^{d_{\textrm{hidden}}}$ is a sequence of hidden states and $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ are matrices which parameterize the LDS. We assume w.l.o.g. that $\mathbf{D}=0$. The problem of learning in dynamical systems refers to predicting the next observation $\mathbf{y}_{t}$ given all previous inputs $\mathbf{u}_{1:t}$.
We consider the problem of learning linear dynamical systems in the online learning setting. In this setting, we measure performance of an algorithm by its regret. Regret is a robust performance measure as it does not assume a specific generative or noise model. It is defined as the difference between the loss incurred by our predictor and that of the best predictor in hindsight belonging to a certain class, i.e.
$\mbox{{Regret}}_{T}({\mathcal{A}})=\sum_{t}\ell(\mathbf{y}_{t},\hat{\mathbf{y}}_{t}^{\mathcal{A}})-\min_{\pi\in\Pi}\sum_{t}\ell(\mathbf{y}_{t},\hat{\mathbf{y}}_{t}^{\pi}).$ (5)
Here ${\mathcal{A}}$ refers to our predictor, $\hat{\mathbf{y}}_{t}^{\mathcal{A}}$ is its prediction at time $t$, $\ell$ is a loss function. Further, $\pi$ is a predictor from the class $\Pi$, for example all predictions that correspond to noiseless linear dynamical systems of the form $\hat{\mathbf{y}}_{t}^{\pi}=\sum_{i=0}^{t}\mathbf{C}\mathbf{A}^{i}\mathbf{B}\mathbf{u}_{t-i}$.
The fundamental problem of learning in linear dynamical systems has been studied for many decades, and we highlight several key approaches below:
  1. 1.
    System identification refers to the method of recovering $\mathbf{A},\mathbf{B},\mathbf{C}$ from the data. This is a hard non-convex problem that is known to fail if the system is marginally stable, i.e. has eigenvalues that are close to one.
  2. 2.
    The (auto) regression method predicts according to $\hat{\mathbf{y}}_{t}=\sum_{i=1}^{h}\mathbf{M}_{i}\mathbf{u}_{t-i}$. The coefficients $\mathbf{M}_{i}$ can be learned using convex regression. The downside of this approach is that if the spectral radius of $\mathbf{A}$ is $1-\delta$, it can be seen that $\sim\frac{1}{\delta}$ terms are needed.
  3. 3.
    The regression method can be further enhanced with “closed loop" components, that regress on prior observations $\mathbf{y}_{t-1:1}$. It can be shown using the Cayley-Hamilton theorem that using this method, $d_{h}$ components are needed to learn the system, where $d_{h}$ is the hidden dimension of $\mathbf{A}$.
  4. 4.
    Filtering involves recovering the state $x_{t}$ from observations. While Kalman filtering is optimal under specific noise conditions, it generally fails in the presence of marginal stability and adversarial noise.
  5. 5.
    Finally, spectral filtering combines the advantages of all methods above. It is an efficient method, its complexity does not depend on the hidden dimension, and works for marginally stable systems. However, spectral filtering requires $\mathbf{A}$ to be symmetric, or diagonalizable under the real numbers.
Our universal sequence preconditioning method gives new results for learning in linear dynamical systems. Algorithms 3 and 4 are efficient methods for general LDS as long as the imaginary components of all eigenvalues is bounded from above, without dependence on the hidden dimension. The condition that the eigenvalues have bounded complex component is in fact necessary in general, as we explain in Section A.1.
Table 1 below summarizes the state of the art in terms of the fundamental problem of learning linear dynamical system, and how our universal sequence preconditioning advances it.
Method Marginally stable $d_{\text{hidden}}$-free asymmetric
Sys-ID $\times$ $\times$ $\checkmark$
Regression, open-loop $\times$ $\checkmark$ $\checkmark$
Regression, closed-loop $\checkmark$ $\times$ $\checkmark$
Spectral filtering $\checkmark$ $\checkmark$ $\times$
[…]
Table 1: Methods for learning LDS: our method (UPS) can learn certain asymetric matrices, whose complex eigenvalues have bounded imaginary components.
4 Analysis Overview
In this section we give the main intuition for universal sequence preconditioning. The crux of our analysis and formal proof of theoretical guarantees are deferred to the appendix.
We start with the main intuition applied to linear dynamical systems. If $(\mathbf{u}_{1:T},\mathbf{y}_{1:T})$ is a linear dynamical system parameterized by $(\mathbf{A},\mathbf{B},\mathbf{C})$
$\mathbf{y}_{t}=\sum_{s=1}^{t}\mathbf{C}\mathbf{A}^{t-s}\mathbf{B}\mathbf{u}_{s}.$
Given coefficients $\mathbf{c}_{0:n}=(\mathbf{c}_{0},\dots,\mathbf{c}_{n})$ let
$p_{n}^{\mathbf{c}}(x)\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i=0}^{n}\mathbf{c}_{i}x^{n-i}.$ (6)
Occasionally we will just write $p_{n}(\cdot)$ when we do not need to specify its coefficients $\mathbf{c}_{0:n}$. Let’s consider the convolution of $\mathbf{y}_{t:t-n}$ with coefficients $\mathbf{c}_{0:n}$ and observe that it yields the following equation
$\sum_{i=0}^{n}c_{i}\mathbf{y}_{t-i}=\sum_{s=0}^{n-1}\sum_{i=0}^{s}c_{i}\mathbf{C}\mathbf{A}^{s-i}\mathbf{B}\mathbf{u}_{t-s}+\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}^{\mathbf{c}}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s}.$
If we take $\mathbf{c}_{0}=1$ (i.e. we have a monic polynomial), we can re-write $\mathbf{y}_{t}$ in terms of several expressions as follows,
$\mathbf{y}_{t}=-\underbrace{\sum_{i=1}^{n}c_{i}\mathbf{y}_{t-i}}_{\aleph_{0}}+\underbrace{\sum_{s=0}^{n-1}\sum_{i=0}^{s}c_{i}\mathbf{C}\mathbf{A}^{s-i}\mathbf{B}\mathbf{u}_{t-s}}_{\aleph_{1}}+\underbrace{\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}^{\mathbf{c}}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s}}_{\aleph_{2}}.$ (7)
This expression highlights our approach as the tension between three terms as follows:
  1. $\aleph_{0}$
    This term is the universal preconditioning part. It depends only on the convolutional coefficients and not on any learning algorithm.
  2. $\aleph_{1}$
    This term can be learned by convex relaxation and then regression over the coefficients, for example by denoting
    $\mathbf{Q}^{\textrm{learned}}_{s}=\sum_{i=0}^{s}c_{i}\mathbf{C}\mathbf{A}^{s-i}\mathbf{B}.$
    The diameter of the coefficient $\mathbf{Q}_{s}$ depends on the convolutional terms $c_{i}$.
  3. $\aleph_{2}$
    This last term has polynomial $p_{n}^{\mathbf{c}}(A)$ which depends on the coefficients $\mathbf{c}_{0:n}$. By a careful choice, we can force this term to be very small, or alternatively we can learn it via spectral filtering.
The main insight we derive from this expression is the inherent tension between two terms $\aleph_{1},\aleph_{2}$. The polynomial $p_{n}^{\mathbf{c}}(x)$ and its corresponding coefficients $c_{0},\dots,c_{n}$ control two terms that are competing:
  1. 1.
    The auto-regressive coefficients grow larger with the degree $n$ of the polynomial and the magnitude of the coefficients $c_{i}$. A higher degree polynomial and larger coefficients increase the diameter of the search space over the auto-regressive coefficients, and therefore increase the regret bound stemming from the $\aleph_{1}$ component learning.
  2. 2.
    On the other hand, a larger search space can allow a broader class of polynomials $p_{n}(\cdot)$ which can better control of the magnitude of $p_{n}(\mathbf{A})$, and therefore reduce the search space of the $\aleph_{2}$ component.
Our two algorithms and theorems differ on their approach to control $\aleph_{2}$, as we now explain.
4.1 Truncation and Convex Regression
Our first, and simpler, Algorithm3 truncates the $t-n$ terms of the expression $\aleph_{2}$. Let $\lambda(\mathbf{A})$ denote the set of eigenvalues of $\mathbf{A}$. By the simple magnitude bound of
$\|\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s}\|~{}\leq~{}\max_{\lambda\in\lambda(\mathbf{A})}|p_{n}(\lambda)|\cdot T\cdot\|\mathbf{C}\|\cdot\|\mathbf{B}\|,$
the error of ignoring this term can be very small if $\max_{\lambda(\mathbf{A})}|p_{n}(\lambda(\mathbf{A}))|$ is small. In the proof of Theorem 2.1 in Appendix B we show that the regret for a generic polynomial $p_{n}^{\mathbf{c}}$ defined by coefficients $\mathbf{c}_{0:n}$ is
$\sum_{t=1}^{T}\|\mathbf{y}_{t}-\hat{\mathbf{y}}_{t}\|_{1}~{}\leq~{}\frac{3}{2}Cn^{2}\sqrt{d_{\textrm{out}}}\|\mathbf{c}\|_{1}\sqrt{T}+C\max_{\lambda\in\mathcal{D}}|p_{n}^{\mathbf{c}}(\lambda)|T^{2},$
where $\mathcal{D}$ is the region where $\mathbf{A}$ is allowed to have eigenvalues (see Theorem B.1). Therefore, to get sublinear regret, we must choose a polynomial which has bounded $\ell_{1}$ norm of its coefficients, while also exhibits very small infinity norm on the domain of $\mathbf{A}$’s eigenvalues.
We note that a simple version of this phenomenon is exhibited by the characteristic polynomial of $\mathbf{A}$. Let $p_{n}(x)$ be the characteristic polynomial for $\mathbf{A}$ and note that it has degree hidden dimension, i.e. $n=d_{h}$. Then by the Cayley-Hamilton theorem, $p_{n}(\mathbf{A})=0$. This means that $\aleph_{2}$ term is canceled out completely. However – the resulting algorithm would require learning hidden dimension many parameters, which is prohibitive.
4.2 Universal Preconditioning and Spectral Filtering
Our second approach, based on spectral filtering, learns the $\aleph_{2}$ term rather than ignoring it. Therefore it has the chance to accumulate much less error, but can take longer to learn. At a very high level, the proof proceeds by exploiting the fact that $p_{n}(\mathbf{A})$ shrinks the size of the learnable domain.
The main guarantee for the spectral filtering algorithm is given in Theorem C.2, which states that Algorithm 4 instantiated with some choice of polynomial $p_{n}^{\mathbf{c}}(\cdot)$ achieves regret
$\tilde{O}\left(\left(n\|\mathbf{c}\|_{1}+T^{7/6}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\right)(n+k)\sqrt{d_{\textrm{out}}}\sqrt{T}\right).$
Both this theorem, as well as our new guarantee for convex regression, leads us to the following question: Is there a universal choice of polynomial $p_{n}(x)$, where $n$ is independent of hidden dimension, which guarantees sublinear regret?
4.3 Using the Chebyshev Polynomial over the Complex Plane
For the real line, the answer to this question is known to be positive using the Chebyshev polynomials of the first kind. In general, the $n^{\textrm{th}}$ (monic) Chebyshev polynomial $M_{n}(x)$ satisfies $\max_{x\in[-1,1]}|M_{n}(x)|~{}\leq~{}2^{-(n-1)}$. However, we are interested in a more general question over the complex plane. Since we care about linear dynamical systems that evolve according to a general asymetric matrix, we need to extending our analysis to $\mathbb{C}_{\beta}$. This is a nontrivial extension since, in general, functions that are bounded on the real line can grow exponentially on the complex plane. Indeed, $2^{n-1}M_{n}(x)=\cos(n\arccos(x))$ and while $\cos(x)$ is bounded within $[-1,1]$ for any $x\in{\mathcal{R}}$, over the complex numbers we have $\cos(z)=\frac{1}{2}(e^{iz}+e^{-iz}),$ which is unbounded. Thus, we analyze the Chebyshev polynomial on the complex plane and provide the following bound.
Lemma 4.1.
Let $z\in\mathbb{C}$ be some complex number with magnitude $|\alpha|~{}\leq~{}1$. Let $M_{n}(\cdot)$ denote the $n$-th monic Chebyshev polynomial. If $|\arg(z)|~{}\leq~{}1/64n^{2}$, then $|M_{n}(z)|~{}\leq~{}1/2^{n-2}$.
We provide the proof in Appendix D. We also must analyze the magnitude of the coefficients of the Chebyshev polynomial, which can grow exponentially with $n$. We provide the following result.
Lemma 4.2.
Let $M_{n}(\cdot)$ have coefficients $c_{0},\dots,c_{n}$. Then $\max_{k=0,\dots,n}|c_{k}|~{}\leq~{}2^{0.3n}$.
The proof of Lemma 4.2 is in Appendix D. Together, these two lemmas are the fundamental building block for universal sequence preconditioning and for obtaining our new regret bounds.
5 Experimental Evaluation
We empirically validate that convolutional preconditioning with Chebyshev or Legendre coefficients yields significant online regret improvements across various learning algorithms and data types. Below we summarize our data generation, algorithm variants, hyperparameter tuning, and evaluation metrics.
5.1 Synthetic Data Generation
We generate $N=200$ sequences of length $T=2000$ via three mechanisms: (i) a noisy linear dynamical system, (ii) a noisy nonlinear dynamical system, and (iii) a noisy deep RNN. Inputs $\mathbf{u}_{1:T}\sim\mathcal{N}(0,I)$.
Linear Dynamical System.
Sample $(\mathbf{A},\mathbf{B},\mathbf{C})$ with $\mathbf{A}\in{\mathcal{R}}^{300\times 300}$ having eigenvalues $\{z_{j}\}$ drawn uniformly in the complex plane subject to $\mathrm{Im}(z_{j})\leq\tau_{\rm thresh}$ and $L\leq|z_{j}|\leq U$, and $\mathbf{B},\mathbf{C}\in{\mathcal{R}}^{300}$. Then
$\mathbf{x}_{t}=\mathbf{A}\,\mathbf{x}_{t-1}+\mathbf{B}\,\mathbf{u}_{t},\quad\mathbf{y}_{t}=\mathbf{C}\,\mathbf{x}_{t}+\epsilon_{t},\;\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}I).$
Nonlinear Dynamical System.
Similarly sample $(\mathbf{A}_{1},\mathbf{B}_{1},\mathbf{C})$ and $(\mathbf{A}_{2},\mathbf{B}_{2})$ with $\mathbf{A}_{i}\in{\mathcal{R}}^{10\times 10}$, $\mathbf{B}_{i},\mathbf{C}\in{\mathcal{R}}^{10}$. Then
$\mathbf{x}_{t}^{(0)}=\mathbf{A}_{1}\mathbf{x}_{t-1}+\mathbf{B}_{1}\mathbf{u}_{t},\;\mathbf{x}_{t}^{(1)}=\sigma\bigl{(}\mathbf{x}_{t}^{(0)}\bigr{)},\;\mathbf{x}_{t}=A_{2}\mathbf{x}_{t}^{(1)}+B_{2}\mathbf{u}_{t},\;\mathbf{y}_{t}=\mathbf{C}\mathbf{x}_{t}+\epsilon_{t}.$
Deep RNN.
We randomly initialize a sparse 10-layer stack of LSTMs with hidden dimension $100$ and ReLU nonlinear activations. Given $\mathbf{u}_{1:T}$ we use this network to generate $\mathbf{y}_{1:T}$.
5.2 Algorithms and Preconditioning Variants
We evaluate the following methods: (1) Regression (Alg. 3) , (2) Spectral Filtering (Alg. 4) , (3) DNN Predictor: $n$-layer LSTM with dims $[d_{1},\dots,d_{n}]$, ReLU.
Each method is applied with one of:
  1. 1.
    Baseline: no preconditioning
  2. 2.
    Chebyshev: $\mathbf{c}_{0:n}$ are the coefficients for the $n$th-Chebyshev polynomial
  3. 3.
    Legendre: $\mathbf{c}_{0:n}$ are the coefficients for the $n$th-Legendre polynomial
  4. 4.
    Learned: $\mathbf{c}_{0:n}$ is a parameter learned jointly with the model parameters
We test polynomial degrees $n\in\{2,5,10,20\}$. This choice of degrees shows a rough picture of the impact of $n$. For Chebyshev and Legendre, once the degree is higher than $5-10$ the performance degrades since $\|\mathbf{c}\|_{1}$ gets very large (see our Lemma 4.2 which shows that these coefficients grow exponentially fast). For the learned coefficients we find that the performance gains are minimal past degree 20, as can be seen in Tables 23, and 4.
Hyperparameter Tuning.
To ensure fair comparison, for each algorithm and conditioning $\mathbf{c}$ variant we perform a grid search over learning rates $\eta\in\{10^{-3},10^{-2},10^{-1}\}$, selecting the one minimizing average regret across the $N$ sequences. In the case of the learned coefficients, we sweep over the 9 pairs of learning rates $(\eta_{\textrm{model}},\eta_{\textrm{coefficients}})\in\{10^{-3},10^{-2},10^{-1}\}\times\{10^{-3},10^{-2},10^{-1}\}$.
5.3 Results
Tables 24 report the mean $\pm$ std of the absolute error over the final 200 predictions, averaged across 200 runs. In the linear and nonlinear cases we train a 2-layer DNN (dims $(64,128)$); for RNN-generated data we match the 10-layer (100-dim) generator.
Key observations:
  • Preconditioning drastically reduces baseline errors for all algorithms and data types.
  • Chebyshev and Legendre yield nearly identical gains.
  • Improvements decay as the complex threshold $\tau_{\rm thresh}$ increases, consistent with our theoretical results which must bound $\mathrm{Im}(z_{j})$.
  • Learned coefficients excel with regression and spectral filtering but destabilize the DNN on nonlinear and RNN-generated data.
Setting Baseline Chebyshev Legendre Learned Deg. 2 Deg. 5 Deg. 10 Deg. 2 Deg. 5 Deg. 10 Deg. 2 Deg. 5 Deg. 10 Deg. 20 Regression $\tau_{\rm thresh}=0.01$ 0.74 $\pm$ 0.28 0.25 $\pm$ 0.09 0.15 $\pm$ 0.07 0.77 $\pm$ 0.31 0.36 $\pm$ 0.13 0.14 $\pm$ 0.06 0.64 $\pm$ 0.26 0.52 $\pm$ 0.19 0.27 $\pm$ 0.11 0.17 $\pm$ 0.07 0.24 $\pm$ 0.09 $\tau_{\rm thresh}=0.1$ 1.92 $\pm$ 0.81 0.84 $\pm$ 0.27 0.66 $\pm$ 0.18 1.90 $\pm$ 0.67 1.10 $\pm$ 0.40 0.63 $\pm$ 0.17 1.66 $\pm$ 0.58 1.34 $\pm$ 0.43 0.57 $\pm$ 0.14 0.55 $\pm$ 0.14 0.56 $\pm$ 0.14 s $\tau_{\rm thresh}=0.9$ 2.47 $\pm$ 0.89 1.59 $\pm$ 0.56 2.18 $\pm$ 0.79 2.68 $\pm$ 0.48 1.64 $\pm$ 0.58 1.94 $\pm$ 0.70 2.63 $\pm$ 0.45 1.73 $\pm$ 0.59 0.83 $\pm$ 0.32 0.68 $\pm$ 0.27 0.63 $\pm$ 0.26 Spectral Filtering $\tau_{\rm thresh}=0.01$ 5.94 $\pm$ 3.37 1.72 $\pm$ 0.95 0.69 $\pm$ 0.38 3.25 $\pm$ 1.79 2.78 $\pm$ 1.56 0.66$\pm$ 0.36 2.74 $\pm$ 1.51 1.99 $\pm$ 0.94 0.54$\pm$ 0.29 0.55 $\pm$ 0.30 0.61 $\pm$ 0.32 $\tau_{\rm thresh}=0.1$ 0.89 $\pm$ 0.34 0.42 $\pm$ 0.11 0.34 $\pm$ 0.07 0.86 $\pm$ 0.28 0.54 $\pm$ 0.17 0.33 $\pm$ 0.07 0.76 $\pm$ 0.24 0.69 $\pm$ 0.28 0.31 $\pm$ 0.06 0.37 $\pm$ 0.08 0.45 $\pm$ 0.28 $\tau_{\rm thresh}=0.9$ 10.17 $\pm$ 8.80 9.87 $\pm$ 8.90 12.66 $\pm$ 8.18 32.90 $\pm$ 22.02 9.42 $\pm$ 8.39 11.53 $\pm$ 7.46 28.83 $\pm$ 19.24 7.93 $\pm$ 4.42 6.31 $\pm$ 4.19 5.73 $\pm$ 3.80 5.86 $\pm$ 4.02 2-layer DNN $\tau_{\rm thresh}=0.01$ 4.49 $\pm$ 2.02 2.31 $\pm$ 1.18 2.62 $\pm$ 1.52 10.36$\pm$ 6.05 2.79 $\pm$ 1.34 2.35 $\pm$ 1.35 8.92 $\pm$ 5.20 2.89 $\pm$ 1.24 1.56 $\pm$ 0.64 0.79 $\pm$ 0.32 0.40 $\pm$ 0.16 $\tau_{\rm thresh}=0.1$ 9.41 $\pm$ 7.34 2.66 $\pm$ 1.76 1.52 $\pm$ 0.72 4.65 $\pm$ 3.03 4.24$\pm$ 3.06 1.44 $\pm$ 0.71 4.03 $\pm$ 2.64 6.54 $\pm$ 4.32 3.22 $\pm$ 1.93 1.59 $\pm$ 0.95 0.80$\pm$ 0.48 $\tau_{\rm thresh}=0.9$ 2.45 $\pm$ 1.31 2.24 $\pm$ 1.34 3.49 $\pm$ 2.27 11.48 $\pm$ 8.15 2.17 $\pm$ 1.25 3.10 $\pm$ 2.00 9.97 $\pm$ 7.05 1.29 $\pm$ 0.66 0.71 $\pm$ 0.34 0.43 $\pm$ 0.18 0.24 $\pm$ 0.11
Table 2: Linear dynamical system data (detailed in Sec. 5.1) across varying complex threshold $\tau_{\rm thres}$.
Setting Baseline Chebyshev Legendre Learned Deg. 2 Deg. 5 Deg. 10 Deg. 2 Deg. 5 Deg. 10 Deg. 2 Deg. 5 Deg. 10 Deg. 20 Spectral Filtering $\tau_{\rm thresh}=0.01$ 153.8 $\pm$ 15.7 43.7$\pm$ 19.4 0.92$\pm$ 0.31 0.26$\pm$ 0.15 78.4$\pm$ 34.7 2.82$\pm$ 1.17 0.34$\pm$ 0.19 1.04$\pm$ 0.29 0.07$\pm$ 0.03 0.05$\pm$ 0.03 0.07$\pm$ 0.03 $\tau_{\rm thresh}=0.01$ 124.1$\pm$ 68.4 33.4$\pm$ 17.3 2.02$\pm$ 1.23 0.36$\pm$ 0.31 56.7$\pm$ 30.5 3.81$\pm$ 1.67 0.35$\pm$ 0.26 2.85$\pm$ 1.13 0.04$\pm$ 0.01 0.02$\pm$ 0.01 0.07$\pm$ 0.02 $\tau_{\rm thresh}=0.9$ 165.5$\pm$ 84.5 43.41$\pm$ 16.39 1.27$\pm$ 0.37 1.90$\pm$ 0.80 76.9$\pm$ 29.1 3.32$\pm$ 1.09 1.64$\pm$ 0.69 1.79$\pm$ 0.61 0.10$\pm$ 0.04 0.12$\pm$ 0.05 0.17$\pm$ 0.07 2-layer DNN $\tau_{\rm thresh}=0.01$ 10.46$\pm$ 5.10 3.45$\pm$ 1.19 0.19$\pm$ 0.16 0.43$\pm$ 0.22 3.55$\pm$ 1.56 0.18$\pm$ 0.14 0.37$\pm$ 0.19 45.40$\pm$ 11.35 25.34$\pm$ 9.06 12.73$\pm$ 4.60 6.43$\pm$ 2.32 $\tau_{\rm thresh}=0.1$ 4.20$\pm$ 1.10 3.38$\pm$ 0.79 0.14$\pm$ 0.05 0.41$\pm$ 0.16 3.42$\pm$ 0.79 0.30$\pm$ 0.14 0.36$\pm$ 0.13 67.70$\pm$ 28.40 31.99$\pm$ 11.78 15.95$\pm$ 5.87 8.01$\pm$ 2.95 $\tau_{\rm thresh}=0.9$ 6.72$\pm$ 2.35 2.45$\pm$ 1.12 0.08$\pm$ 0.03 0.28$\pm$ 0.15 3.35$\pm$ 0.98 0.20$\pm$ 0.07 0.24$\pm$ 0.13 58.65$\pm$ 17.09 28.74$\pm$ 8.07 14.39$\pm$ 4.06 7.26$\pm$ 2.05
Table 3: Nonlinear data (detailed in Sec. 5.1) across varying complex threshold $\tau_{\rm thres}$.
Setting Baseline Chebyshev Legendre Learned Deg. 2 Deg. 5 Deg. 10 Deg. 2 Deg. 5 Deg. 10 Deg. 2 Deg. 5 Deg. 10 Deg. 20 10-layer DNN 0.54 $\pm$ 0.23 0.29 $\pm$ 0.10 0.08$\pm$ 0.03 0.13 $\pm$ 0.05 0.37 $\pm$ 0.14 0.09$\pm$ 0.03 0.12$\pm$ 0.04 1.49$\pm$ 0.93 2.13$\pm$ 1.05 1.04$\pm$ 0.51 0.5 $\pm$ 0.24
Table 4: Performance (average absolute error of the last 200 predictions) of a 10-layer DNN (detailed in Sec. 5.2) on data generated from the same model (detailed in Sec. 5.1).
6 Conclusion and Discussion
6.1 Why should preconditioning be universal?
There are many settings in machine learning where universal, rather than learned, rules have proven very efficient. For example, physical laws of motion can be learned directly from observation data. However, Newton’s laws of motion succinctly crystallize very general phenomenon, and have proven very useful for large scale physics simulation engines.
Similarly, in the theory of mathematical optimization, adaptive gradient methods have revolutionized deep learning. Their derivation as a consequence of regularization in online regret minimization is particularly simple duchi2011adaptive , and thousands of research papers have not dramatically improved the initial basic ideas. These optimizers are, at the very least, a great way to initialize learned optimizers wichrowska2017learned .
By analogy, our thesis in this paper is that universal preconditioning based on the solid theory of dynamical systems can be applicable to many domains or, at the very least, an initialization for other learning methods.
6.2 Summary
This paper concerns sequence preconditioning, a topic that was initially studied in the context of statistical time series analysis. We expand the simple techniques of differencing, and devise a universal preconditioning method based on orthogonal polynomials.
We apply our method to learning in linear dynamical systems, a fundamental setting in learning and control. We have significantly expanded the known systems that can be efficiently learned to those that are:
  • Marginally stable: systems whose their spectral radius is arbitrarily close to one. Intuitively this captures long term effects.
  • Large hidden dimension: systems with larger hidden dimension are more expressive.
  • General transition matrices: systems that can have asymmetric transition matrices.
To our knowledge, no previous work is able to provably learn systems that have this combination without depending on hidden dimension. We are able to accomplish this by introducing new techniques: universal preconditioning with the Chebychev polynomial coefficients.
We further improve performance bounds by introducing a new method of spectral filtering, extending it to the complex domain under mild assumptions on the boundedness of the imaginary domain of the system eigenstructure.
Appendix A Related work
Preconditioning for sequence prediction.
Preconditioning in the context of time series analysis has roots in the classical work of Box and Jenkins [13]. In this foundational text they propose differencing as a method for making the time series stationary, and thus amenable to statistical learning techniques such as ARMA (auto-regressive moving average) [8]. The differencing operator can be applied numerous times, and for different lags, giving rise to the ARIMA family of forecasting models.
Identifying the order of an ARIMA model, and in particular the types of differencing needed to make a series stationary, is a hard problem This is a special case of the problem we consider: differencing corresponds to certain coefficients of preconditioning the time series, whereas we consider arbitrary coefficients.
Background on control of linear dynamical systems.
Linear dynamical systems are the most fundamental and basic model in control theory, and have been studied for more than a century. For a thorough introduction, see the texts [12, 32, 24].
A rigorous proof that the Caley-Hamilton theorem implies that $d_{h}$ learned closed-loop components are sufficient to learn any LDS is given in [7, 23].
The seminal work of Kalman on state-space representation and filtering [26] allows to learn any LDS with hidden-dimension parameters under stochastic and generative assumptions. Closed loop auto-regressive learning subsumes Kalman filtering in the presence of adversarial noise, see e.g. [28]. [21] provide a method to learn a marginally stable LDS in the presence of bounded adversarial noise and asymmetric matrices, however their regret bound depends on the hidden dimension of the system. More recently, [10] use tensor methods to learn a LDS with stochastic noise without dependence on the system’s condition number. However, their method also has a complexity that depends on the hidden dimension of the system.
In this work we consider regret in the context of learning linear dynamical systems. This is related to, but different from, control of the systems. We survey regret minimization for control next.
Regret for classical control models.
The first works addressing control in the machine learning community assume either no perturbation in the dynamics at all, or i.i.d. Gaussian perturbations. Much of this work has considered obtaining low regret in the online LQR setting [2, 18, 29, 17] where a fully-observed linear dynamic system is drive by i.i.d. Gaussian noise via $x_{t+1}=\mathbf{A}x_{t}+\mathbf{B}\mathbf{u}_{t}+w_{t}$, and the learner incurs constant quadratic state and input cost $\ell(x,u)=\frac{1}{2}x^{\top}\mathbf{Q}x+\frac{1}{2}u^{\top}Ru$. The optimal policy for this setting is well-approximated by a state feedback controller $\mathbf{u}_{t}=K\mathbf{u}_{t}$, where $K$ is the solution to the Discrete Algebraic Ricatti Equation (DARE), and thus regret amounts to competing with this controller. Recent algorithms [29, 17] attain $\sqrt{T}$ regret for this setting, with polynomial runtime and polynomial regret dependence on relevant problem parameters. A parallel line of work by [16] establish $\sqrt{T}$ in a variant of online LQR where the system is known to the learner, noise is stochastic, but an adversary selects quadratic loss functions $\ell_{t}$ at each time $t$. Again, the regret is measured with respect to a best-in-hindsight state feedback controller.
Provable control in the Gaussian noise setting via the policy gradient method was studied in [20]. Other relevant work from the machine learning literature includes tracking adversarial targets [1].
The non-stochastic control problem.
The most accepted and influential control model stemming from machine learning community was established in [4], who obtain $\sqrt{T}$-regret in the more general and challenging setting where the Lipschitz loss function and the perturbations are adversarially chosen. The key insight behind this result is combining an improper controller parametrization know as disturbance-based control with online convex optimization with memory due to [9]. Follow up work by [6] achieves logarithmic pseudo-regret for strongly convex, adversarially selected losses and well-conditioned stochastic noise. Further extensions were made for linear control with partial observation [30], system identification with adversarial noise [15], and many more settings surveyed in [24].
Online learning and online convex optimization.
We make extensive use of techniques from the field of online learning and regret minimization in games [14, 22]. Following previous work from machine learning, we consider regret minimization in sequence prediction, where the underlying sequence follows a linear dynamical system.
Spectral filtering for learning linear dynamical systems.
The spectral filtering technique was put forth in [25] for learning symmetric linear dynamical systems. In [23], the technique was extended for more general dynamical systems using closed loop regression, however this required hidden-dimension parameters and polynomial dependence on the approximation guarantee. Spectral filtering techniques were recently used in non-linear sequence prediction, notably in the context of large language models, albeit with no theoretical guarantees [7]. As convolutional models, these methods are attractive for sequence prediction due to faster generation and inference as compared to attention-based models [5].
In this paper we dramatically improve the spectral filtering technique and broaden its applicability in two major aspects: First, for general asymetric linear dynamical systems we remove the dependence on the hidden dimension. Second, we improve the dependence of the number of learned parameters from polynomial to logarithmic.
A.1 Preconditioning for Linear Dynamical Systems
Linear dynamical systems (LDS) are perhaps the most basic and well-studied dynamical systems in engineering and control science. They form a fundamental class of models used in control theory, time-series forecasting, and machine learning. Given input vectors $\mathbf{u}_{1},\dots,\mathbf{u}_{T}\in{\mathcal{R}}^{d_{\textrm{in}}}$, the system generates a sequence of output vectors $\mathbf{y}_{1},\dots\mathbf{y}_{T}\in{\mathcal{R}}^{d_{\textrm{out}}}$ according to the law
$\mathbf{x}_{t+1}$ $=\mathbf{A}\mathbf{x}_{t}+\mathbf{B}\mathbf{u}_{t}$
$\mathbf{y}_{t}$ $=\mathbf{C}\mathbf{x}_{t}+\mathbf{D}\mathbf{u}_{t},$
where $\mathbf{x}_{0},\dots,\mathbf{x}_{T}\in{\mathcal{R}}^{d_{\textrm{hidden}}}$ is a sequence of hidden states with hidden dimension $d_{\textrm{hidden}}$, $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ are matrices which parameterize the LDS, and we ignore dynamics and observation noise for clarity. The problem of learning a partially observable LDS refers to predicting the next observation $\mathbf{y}_{t}$ given all previous inputs $\mathbf{u}_{1:t}$ and outputs $\mathbf{y}_{1:{t-1}}$. This problem remains a central challenge, particularly when the system is marginally stable, i.e. its transition matrix $\mathbf{A}$ has spectral radius close to one. Until now, there has been been no known efficient method to learn a general, marginally stable LDS in an online setting without dependence on hidden dimension $d_{\textrm{hidden}}$.
While several methods exist that can learn in the presence of asymmetric transition matrices [26, 10, 21], their performance depends on hidden dimension. On the other hand, spectral filtering methods [25] achieve regret which is independent of hidden dimension, even for marginally stable systems. However, these spectral filtering methods were limited to systems with symmetric transition matrices. In contrast, real-world dynamical systems often have asymmetric transition matrices with large hidden dimension, necessitating a more general approach. In this paper, we provide such an approach by extending the theory of spectral filtering to handle asymmetric systems, as long as the complex component of their eigenvalues is bounded.
A.2 Memory Capacity of Linear Dynamical Systems
The hidden dimension $d_{h}$, which is the dimension of the transition matrix $\mathbf{A}$, plays a significant role in the expressive power of LDS. One of the most important features of the hidden dimension is that an LDS can memorize and recall inputs from up to ${d_{\textrm{hidden}}}$ iterations in the past.
This can be seen with the system where $\mathbf{B},\mathbf{C}$ are identity, and $\mathbf{A}$ is given by the permutation matrix
$\mathbf{A}_{d_{\textrm{hidden}}}^{\mathrm{perm}}=\begin{bmatrix}0&0&\cdots&0&1\\ 1&0&\cdots&0&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&0&0\\ 0&0&\cdots&1&0\\ \end{bmatrix},$
which implements the memory operator $\mathbf{y}_{t}=\mathbf{u}_{t-d_{h}}$. Observe that any method which uses fewer than $d_{\textrm{hidden}}$ parameters will fail to implement this memory operator and therefore, for general linear dynamical systems, $d_{\textrm{hidden}}$ parameters are necessary.
Seemingly, this contradicts our promised results, which allows for learning a general LDS without hidden dimension dependence. The explanation is in the spectrum of the system. Notice that the eigenvalues of the permutation matrix $\mathbf{A}$ above are the $d_{\textrm{hidden}}$ roots of unity given by
$\lambda_{1},...,\lambda_{d_{\textrm{hidden}}}\in\left\{e^{2\pi i\frac{k}{d}},k=1,2,...,d_{\textrm{hidden}}\right\}.$
Note that these eigenvalues have complex component as large as $1-1/d_{\textrm{hidden}}$.
Although in general a LDS can express signals with $d_{\textrm{hidden}}$ memory, and thus intuitively might require $d_{\textrm{hidden}}$ parameters, there are notable special cases that allow for efficient learning, i.e. learning the LDS with far fewer parameters. A notable case is that of spectral filtering, which allows efficient learning of a symmetric LDS with poly-logarithmic (in the desired accuracy $\varepsilon$) number of parameters. The significance of a symmetric transition matrix $\mathbf{A}$ is that it can be diagonalized over the real numbers.
The natural question that arises is which asymmetric matrices can be learned by spectral filtering efficiently, and which characterization of their spectrum allows for efficient learning?
The answer we offer is surprisingly broad. For a LDS with transition matrix $\mathbf{A}$, let $\lambda_{1},...,\lambda_{d}$ be its complex eigenvalues. We show that we can learn up to $\varepsilon$ accuracy any LDS for which the largest eigenvalue has imaginary part bounded by $\frac{1}{\mathrm{poly}\log\frac{1}{\varepsilon}}$. We note that the spectral radius can be arbitrarily close to, or even equal to, one. The only restriction is on the complex part, which is mildly constrained as a logarithmic function of $\varepsilon$.
As per the permutation matrix example, this dependence is necessary and nearly tight - if the imaginary component of the eigenvalues of $\mathbf{A}$ becomes large, any learning method requires parameterization that depends on the hidden dimension of the system.
Appendix B Proof of Convolutional Preconditioned Regression Performance Theorem 2.1
We will prove Theorem 2.1 first by proving the general result of Algorithm 3 for any choice of preconditioning coefficients $\mathbf{c}_{0:n}$. Then we will apply the Chebyshev coefficients to the result to get Theorem 2.1. For convenience, we restate Theorem 2.1 here.
Theorem B.1 (General Form of Theorem 2.1).
Let $\left\{\mathbf{u}_{t}\right\}_{t=1}^{T}\in{\mathcal{R}}^{d_{\textrm{in}}}$ be any sequence of inputs which satisfy $\|\mathbf{u}_{t}\|_{2}~{}\leq~{}1$ and let $\left\{\mathbf{y}_{t}\right\}_{t=1}^{T}$ be the corresponding output coming from some linear dynamical system $(\mathbf{A},\mathbf{B},\mathbf{C})$ as defined per Eq. 3. Assume that $\|\mathbf{B}\|\|\mathbf{C}\|~{}\leq~{}C_{\textrm{domain}}$. Then the predictions $\hat{\mathbf{y}}_{1},\dots,\hat{\mathbf{y}}_{T}$ from Algorithm 3 satisfy
$\sum_{t=1}^{T}\|\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\|_{1}~{}\leq~{}C_{\textrm{domain}}\left(\frac{3}{2}n^{2}\sqrt{d_{\textrm{out}}}\|\mathbf{c}\|_{1}\sqrt{T}+\max_{\lambda\in\lambda(\mathbf{A})}|p_{n}^{\mathbf{c}}(\lambda)|T^{2}\right).$
Proof of Theorem B.1.
For the remainder of the proof we will denote
$\hat{\mathbf{y}}_{t}(\mathbf{Q})=-\sum_{i=1}^{n}\mathbf{c}_{i}\mathbf{y}_{t-i}+\sum_{j=0}^{n}\mathbf{Q}_{j}\mathbf{u}_{t-j},$
so that Algorithm 3 outputs $\hat{\mathbf{y}}_{t}=\hat{\mathbf{y}}_{t}(\mathbf{Q}^{t})$. Recall we define the domain
$\mathcal{K}=\left\{\left(\mathbf{Q}_{0},\dots,\mathbf{Q}_{n-1}\right)\textrm{ s.t. }\|\mathbf{Q}_{j}\|~{}\leq~{}C_{\textrm{domain}}n\|\mathbf{c}\|_{1}\right\}.$
For convenience, we will use the shorthand $\mathbf{Q}$ to refer to $\left(\mathbf{Q}_{0},\dots,\mathbf{Q}_{n-1}\right)$. First we prove that the regret of Algorithm 3 as compared to the best $\mathbf{Q}^{*}\in\mathcal{K}$ is hindsight is small. Then we prove that the best $\mathbf{Q}^{*}\in\mathcal{K}$ in hindsight achieves small prediction error. Let
$G=\max_{t\in[T]}\|\nabla\mkern-2.5mu_{\mathbf{Q}}\ell_{t}(\mathbf{Q}^{t})\|,$
and let
$D=\max_{\mathbf{Q}^{1},\mathbf{Q}^{2}\in\mathcal{K}}\|\mathbf{Q}^{1}-\mathbf{Q}^{2}\|.$
By Theorem 3.1 from [22],
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t})-\min_{\mathbf{Q}^{*}\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{*})~{}\leq~{}\frac{3}{2}GD\sqrt{T}.$
Therefore it remains to bound $D$ and $G$. First we bound $D$. By definition of $\mathcal{K}$,
$D$ $=\max_{\mathbf{Q}^{1},\mathbf{Q}^{2}\in\mathcal{K}}\|\mathbf{Q}^{1}-\mathbf{Q}^{2}\|$
$=\max_{(\mathbf{Q}^{1}_{0},\dots,\mathbf{Q}^{1}_{n-1}),(\mathbf{Q}^{2}_{0},\dots,\mathbf{Q}^{2}_{n-1})\in\mathcal{K}}\|(\mathbf{Q}^{1}_{0},\dots,\mathbf{Q}^{1}_{n-1})-(\mathbf{Q}^{2}_{0},\dots,\mathbf{Q}^{2}_{n-1})\|$
$~{}\leq~{}\sum_{j=0}^{n-1}\|\mathbf{Q}^{1}_{j}-\mathbf{Q}^{2}_{j}\|$
$~{}\leq~{}2C_{\textrm{domain}}n^{2}\|\mathbf{c}\|_{1}.$
Next we bound the gradient. Recall that
$\ell_{t}(\mathbf{Q})$ $=\|\hat{\mathbf{y}}_{t}(\mathbf{Q})-\mathbf{y}_{t}\|_{1}$
$=\|-\sum_{i=1}^{n}\mathbf{c}_{i}\mathbf{y}_{t-i}+\sum_{j=0}^{n}\mathbf{Q}_{j}\mathbf{u}_{t-j}-\mathbf{y}_{t}\|_{1}.$
Note that, in general, $\nabla\mkern-2.5mu_{\mathbf{A}}\|\mathbf{A}\mathbf{x}-\mathbf{b}\|_{1}=\mathrm{sign}(\mathbf{A}\mathbf{x}-\mathbf{b})\mathbf{x}^{\top}$. Since $\|\mathbf{x}\mathbf{y}^{\top}\|_{F}~{}\leq~{}\|\mathbf{x}\|_{2}\|\mathbf{y}\|_{2}$ we have
$\|\nabla\mkern-2.5mu_{\mathbf{A}}f(\mathbf{A})\|_{F}~{}\leq~{}\sqrt{d}\|\mathbf{x}\|_{2},$
where $d$ is the dimension of $\mathbf{b}$. Using this and the assumption that for any $t\in[T]$, $\|\mathbf{u}_{t}\|_{2}~{}\leq~{}1$, we have
$\|\nabla\mkern-2.5mu_{\mathbf{Q}_{i}}\ell_{t}(\mathbf{Q})\|_{F}$ $~{}\leq~{}\sqrt{d_{\textrm{out}}}\|\mathbf{u}_{t-i}\|_{2}~{}\leq~{}\sqrt{d_{\textrm{out}}}.$
Therefore,
$G=\max_{t\in[T]}\|\nabla\mkern-2.5mu_{\mathbf{Q}}\ell_{t}(\mathbf{Q}^{t})\|_{F}~{}\leq~{}n\sqrt{d_{\textrm{out}}}.$
Thus we have a final regret bound
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t})-\min_{\mathbf{Q}^{*}\in\mathcal{K}}\ell_{t}(\mathbf{Q}^{*})~{}\leq~{}\frac{3}{2}C_{\textrm{domain}}n^{2}\sqrt{d_{\textrm{out}}}\|\mathbf{c}\|_{1}\sqrt{T}.$
Next we show that if $(\mathbf{u}_{1:T},\mathbf{y}_{1:T})$ is a linear dynamical system parameterized by $(\mathbf{A},\mathbf{B},\mathbf{C})$, then for any $t\in[T]$,
$\min_{\mathbf{Q}^{*}\in\mathcal{K}}\ell_{t}(\mathbf{Q}^{*})~{}\leq~{}\|\mathbf{C}\|\|\mathbf{B}\|\cdot\max_{\lambda\in\lambda(\mathbf{A})}|p_{n}^{\mathbf{c}}(\lambda)|\cdot T,$
where $p_{n}^{\mathbf{c}}$ denotes the polynomial
$p_{n}^{\mathbf{c}}(x)\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i=0}^{n}\mathbf{c}_{i}x^{n-i},$
and $\lambda(\mathbf{A})$ denotes the set of eigenvalues of $\mathbf{A}$. Indeed, if $(\mathbf{u}_{1:T},\mathbf{y}_{1:T})$ is a linear dynamical system parameterized by $(\mathbf{A},\mathbf{B},\mathbf{C})$ then
$\mathbf{y}_{t}=\sum_{s=1}^{t}\mathbf{C}\mathbf{A}^{t-s}\mathbf{B}\mathbf{u}_{s}.$
With some linear algebra we get that convolving $\mathbf{y}_{t:t-n}$ with coefficients $\mathbf{c}_{0:n}$ results in
$\sum_{i=0}^{n}\mathbf{c}_{i}\mathbf{y}_{t-i}=\sum_{s=0}^{n-1}\sum_{i=0}^{s}\mathbf{c}_{i}\mathbf{C}\mathbf{A}^{s-i}\mathbf{B}\mathbf{u}_{t-s}+\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}^{\mathbf{c}}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s},$
or equivalently (since $\mathbf{c}_{0}=1$ due to the assertion in Algorithm 3),
$\mathbf{y}_{t}=-\sum_{i=1}^{n}\mathbf{c}_{i}\mathbf{y}_{t-i}+\sum_{s=0}^{n-1}\sum_{i=0}^{s}\mathbf{c}_{i}\mathbf{C}\mathbf{A}^{s-i}\mathbf{B}\mathbf{u}_{t-s}+\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}^{\mathbf{c}}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s}.$
Set $\hat{\mathbf{Q}}_{s}=\sum_{i=0}^{s}\mathbf{c}_{i}\mathbf{C}\mathbf{A}^{s-i}\mathbf{B}$ and set $\hat{\mathbf{Q}}=\left(\hat{\mathbf{Q}}_{0},\dots,\hat{\mathbf{Q}}_{n-1}\right)$. Since we assumed $C_{\textrm{domain}}~{}\geq~{}\|\mathbf{C}\|\|\mathbf{B}\|$,
$\|\hat{\mathbf{Q}}_{i}\|~{}\leq~{}\sum_{j=0}^{i}|\mathbf{c}_{j}|\|\mathbf{C}\mathbf{A}^{i-j}\mathbf{B}\|~{}\leq~{}\|\mathbf{C}\|\|\mathbf{B}\|\sum_{j=0}^{i}|\mathbf{c}_{j}|~{}\leq~{}C_{\textrm{domain}}n\|\mathbf{c}\|_{1}.$
Therefore $\hat{\mathbf{Q}}\in\mathcal{K}$. Moreover,
$\hat{\mathbf{y}}_{t}(\hat{\mathbf{Q}})-\mathbf{y}_{t}$ $=\left(-\sum_{i=1}^{n}\mathbf{c}_{i}\mathbf{y}_{t-i}+\sum_{j=0}^{n-1}\hat{\mathbf{Q}}_{j}\mathbf{u}_{t-j}\right)$
$\qquad-\left(-\sum_{i=1}^{n}\mathbf{c}_{i}\mathbf{y}_{t-i}+\sum_{s=0}^{n-1}\sum_{i=0}^{s}\mathbf{c}_{i}\mathbf{C}\mathbf{A}^{s-i}\mathbf{B}\mathbf{u}_{t-s}+\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}^{\mathbf{c}}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s}\right)$
$=\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}^{\mathbf{c}}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s}.$
Therefore, $\|\hat{\mathbf{y}}_{t}(\hat{\mathbf{Q}})-\mathbf{y}_{t}\|_{1}=\|\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}^{\mathbf{c}}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s}\|_{1}$. Let $\mathbf{A}$ be diagonalized by some $\mathbf{P}$ so that
$\mathbf{A}=\mathbf{P}\mathbf{D}\mathbf{P}^{-1},$
where $\mathbf{D}$ is the diagonalization of $\mathbf{A}$, note that we can assume this w.l.o.g. since the set of dagonalizable matrices over the complex numbers is dense and therefore if $\mathbf{A}$ is not diagonalizable we may apply an aribtrarily small perturbation to it. Then since $\max_{j\in[d_{h}]}|\lambda_{j}(\mathbf{A})|~{}\leq~{}1$,
$\|\mathbf{C}p_{n}^{\mathbf{c}}(\mathbf{A})\mathbf{A}^{j}\mathbf{B}\|$ $=\|\mathbf{C}\mathbf{P}p_{n}^{\mathbf{c}}(\mathbf{D})\mathbf{D}^{j}\mathbf{P}^{-1}\mathbf{B}\|$
$~{}\leq~{}\max_{k\in[d_{h}]}|p_{n}^{\mathbf{c}}(\mathbf{D}_{kk}|\cdot\|\mathbf{C}\mathbf{P}\mathbf{D}^{j}\mathbf{P}^{-1}\mathbf{B}\|$
$=\max_{\lambda\in\lambda(\mathbf{A})}|p_{n}^{\mathbf{c}}(\lambda)|\cdot\|\mathbf{C}\mathbf{A}^{j}\mathbf{B}\|$
$~{}\leq~{}\max_{\lambda\in\lambda(\mathbf{A})}|p_{n}^{\mathbf{c}}(\lambda)|\cdot\|\mathbf{C}\|\|\mathbf{B}\|.$
Thus,
$\|\hat{\mathbf{y}}_{t}(\hat{\mathbf{Q}})-\mathbf{y}_{t}\|_{1}~{}\leq~{}\|\mathbf{C}\|\|\mathbf{B}\|\cdot\max_{\lambda\in\lambda(\mathbf{A})}|p_{n}^{\mathbf{c}}(\lambda)|\cdot T,$
and so (recalling that we showed $\hat{\mathbf{Q}}\in\mathcal{K}$),
$\min_{\mathbf{Q}^{*}\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{*})$ $~{}\leq~{}\sum_{t=1}^{T}\ell_{t}(\hat{\mathbf{Q}})$
$=\sum_{t=1}^{T}\|\hat{\mathbf{y}}_{t}(\hat{\mathbf{Q}})-\mathbf{y}_{t}\|_{1}$
$~{}\leq~{}\|\mathbf{C}\|\|\mathbf{B}\|\cdot\max_{\lambda\in\lambda(\mathbf{A})}|p_{n}^{\mathbf{c}}(\lambda)|\cdot T^{2}.$
Therefore we conclude that as long as $C_{\textrm{domain}}~{}\geq~{}\|\mathbf{C}\|\cdot\|\mathbf{B}\|$,
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t})~{}\leq~{}C_{\textrm{domain}}\left(\frac{3}{2}n^{2}\sqrt{d_{\textrm{out}}}\|\mathbf{c}\|_{1}\sqrt{T}+\max_{\lambda\in\lambda(\mathbf{A})}|p_{n}^{\mathbf{c}}(\lambda)|T^{2}\right).$
  0.0pt1.5ex1.5ex
Next we choose $\mathbf{c}$ to be the coefficients of the $n$-th monic Chebyshev polynomial to get the original theorem, Theorem 2.1, restated here for convenience.
Theorem (Restatement of Theorem 2.1).
Let $\left\{\mathbf{u}_{t}\right\}_{t=1}^{T}\in{\mathcal{R}}^{d_{\textrm{in}}}$ be any sequence of inputs which satisfy $\|\mathbf{u}_{t}\|_{2}~{}\leq~{}1$ and let $\left\{\mathbf{y}_{t}\right\}_{t=1}^{T}$ be the corresponding output coming from some linear dynamical system $(\mathbf{A},\mathbf{B},\mathbf{C})$ as defined per Eq. 3. Assume that $\|\mathbf{B}\|\|\mathbf{C}\|~{}\leq~{}C_{\textrm{domain}}$. Let $\lambda_{1},\dots,\lambda_{d_{h}}$ denote the spectrum of $\mathbf{A}$. If
$\max_{j\in[d_{h}]}|\mathrm{Arg}(\lambda_{j})|~{}\leq~{}\left(64\log_{2}\left(\frac{8T^{3/2}}{3\sqrt{d_{\textrm{out}}}}\right)\right)^{-2},$
then the predictions $\hat{\mathbf{y}}_{1},\dots,\hat{\mathbf{y}}_{T}$ from Algorithm 3 where the preconditioning coefficients $\mathbf{c}_{0:n}$ are chosen to be the coefficients of the $n$-th monic Chebyshev polynomial satisfy
$\frac{1}{T}\sum_{t=1}^{T}\|\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\|_{1}~{}\leq~{}\frac{9C_{\textrm{domain}}\sqrt{d_{\textrm{out}}}\log_{2}^{2}(3T)}{T^{2/13}}.$
Proof.
From Theorem B.1 we have
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t})~{}\leq~{}C_{\textrm{domain}}\left(\frac{3}{2}n^{2}\sqrt{d_{\textrm{out}}}\|\mathbf{c}\|_{1}\sqrt{T}+\max_{\lambda\in\lambda(\mathbf{A})}|p_{n}^{\mathbf{c}}(\lambda)|T^{2}\right).$
By Lemma 4.1, if for any eigenvalue $\lambda$ of $\mathbf{A}$, $|\arg(\lambda)|~{}\leq~{}1/64n^{2}$ then
$\max_{\lambda\in\lambda(\mathbf{A})}|p_{n}^{\mathbf{c}}(\lambda)|~{}\leq~{}\frac{1}{2^{n-2}}.$
Moreover, by Lemma 4.2, $\|\mathbf{c}\|_{1}~{}\leq~{}2^{0.3n}$. Thus the Chebyshev-conditioned predictor class satisfies
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t})~{}\leq~{}C_{\textrm{domain}}\left(\frac{3}{2}n^{2}\sqrt{d_{\textrm{out}}}2^{0.3n}\sqrt{T}+2^{-(n-2)}T^{2}\right).$
Picking
$n=\frac{10}{13}\log_{2}\left(\frac{8}{3\sqrt{d_{\textrm{out}}}}T^{3/2}\right),$
we get
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t})$ $~{}\leq~{}2\left(\frac{3}{2}\left(\frac{10}{13}\log_{2}\left(\frac{8}{3\sqrt{d_{\textrm{out}}}}T^{3/2}\right)\right)^{2}\sqrt{d_{\textrm{out}}}\right)^{10/13}4^{3/13}T^{11/13}$
$~{}\leq~{}9C_{\textrm{domain}}\sqrt{d_{\textrm{out}}}\log_{2}(3T)^{2}T^{11/13}.$
Dividing both sides by $T$ we get the stated result. ${1.5ex}{1.5ex}$ 0.0pt1.5ex1.5ex
Appendix C Proof of Convolutional Preconditioned Spectral Filtering Performance Theorem C.2
First, we prove a more general form that holds for any choice of $\mathbf{c}_{0:n}$ and resulting polynomial $p_{n}^{\mathbf{c}}$.
Theorem C.1.
Let $\left\{\mathbf{u}_{t}\right\}_{t=1}^{T}\in{\mathcal{R}}^{d_{\textrm{in}}}$ be any sequence of inputs which satisfy $\|\mathbf{u}_{t}\|_{2}~{}\leq~{}1$ and let $\left\{\mathbf{y}_{t}\right\}_{t=1}^{T}$ be the corresponding output coming from some linear dynamical system $(\mathbf{A},\mathbf{B},\mathbf{C})$ as defined per Eq. 3. If $R_{Q}~{}\geq~{}\|\mathbf{C}\|\cdot\|\mathbf{B}\|\cdot\|\mathbf{c}\|_{1}$, $R_{M}~{}\geq~{}C\log(T)^{1/6}T^{7/6}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\cdot\|\mathbf{C}\mathbf{B}\|$ and $k~{}\geq~{}C\log^{2}\left(T\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\|\mathbf{C}\mathbf{B}\|/(d_{\textrm{out}}\epsilon)\right)$ (where $C>0$ is a universal constant) then the predictions $\hat{\mathbf{y}}_{1},\dots,\hat{\mathbf{y}}_{T}$ from Algorithm 4 satisfy
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})~{}\leq~{}1+C\|\mathbf{C}\|\|\mathbf{B}\|\left(n\|\mathbf{c}\|_{1}+\log^{2}(T)T^{7/6}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\right)(n+k)\sqrt{d_{\textrm{out}}}\sqrt{T}.$
Critically, we observe that the guaranteed regret bound does not depend on the hidden dimension of the dynamics matrix $\mathbf{A}$. While the general version of Algorithm 4 is interesting in its own right, we show that by choosing the coefficients of the algorithm to be the Chebyshev coefficients we obtain sublinear absolute error when the signal comes from a linear dynamical system.
Theorem C.2.
Let $\left\{\mathbf{u}_{t}\right\}_{t=1}^{T}\in{\mathcal{R}}^{d_{\textrm{in}}}$ be any sequence of inputs which satisfy $\|\mathbf{u}_{t}\|_{2}~{}\leq~{}1$ and let $\left\{\mathbf{y}_{t}\right\}_{t=1}^{T}$ be the corresponding output coming from some linear dynamical system $(\mathbf{A},\mathbf{B},\mathbf{C})$ as defined per Eq. 3. If $R_{Q}~{}\geq~{}\|\mathbf{C}\|\|\mathbf{B}\|\|\mathbf{c}\|_{1}$, $R_{M}~{}\geq~{}C\log(T)^{1/6}T^{7/6}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\|\mathbf{C}\mathbf{B}\|$ and $k~{}\geq~{}C\log^{2}\left(T\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\|\mathbf{C}\mathbf{B}\|/(d_{\textrm{out}}\epsilon)\right)$ (where $C>0$ is a universal constant) and
$\max_{j\in[d_{h}]}|\arg(\lambda_{j})|~{}\leq~{}\left(64\log_{2}\left(\frac{8T^{3/2}}{3\sqrt{d_{\textrm{out}}}}\right)\right)^{-2},$
then the predictions $\hat{\mathbf{y}}_{1},\dots,\hat{\mathbf{y}}_{T}$ from Algorithm 4 where the preconditioning coefficients $\mathbf{c}_{0:n}$ are chosen to be the coefficients of the $n$-th monic Chebyshev polynomial satisfy
$\frac{1}{T}\sum_{t=1}^{T}\|\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\|_{1}~{}\leq~{}\tilde{O}\left(\frac{\|\mathbf{C}\|\|\mathbf{B}\|}{T^{3/13}}\right),$
where $\tilde{O}(\cdot)$ hides polylogarithmic factors in $T$.
Proof.
From Theorem B.1 we have
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})~{}\leq~{}1+C\|\mathbf{C}\|\|\mathbf{B}\|\left(n\|\mathbf{c}\|_{1}+\log^{2}(T)T^{7/6}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\right)(n+k)\sqrt{d_{\textrm{out}}}\sqrt{T}.$
By Lemma 4.1, if for any eigenvalue $\lambda$ of $\mathbf{A}$, $|\arg(\lambda)|~{}\leq~{}1/64n^{2}$ then
$\max_{\lambda\in\lambda(\mathbf{A})}|p_{n}^{\mathbf{c}}(\lambda)|~{}\leq~{}\frac{1}{2^{n-2}}.$
Moreover, by Lemma 4.2, $\|\mathbf{c}\|_{1}~{}\leq~{}2^{0.3n}$. Thus the Chebyshev-conditioned predictor class satisfies
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t})~{}\leq~{}1+C\|\mathbf{C}\|\|\mathbf{B}\|\left(n^{2}2^{0.3n}+\log^{2}(T)T^{7/6}2^{-(n-2)}\right)(n+k)\sqrt{d_{\textrm{out}}}\sqrt{T}.$
Picking
$n=\frac{10}{13}\log_{2}\left(4\log^{2}(T)T^{7/6}\right),$
we get
$n^{2}2^{0.3n}+\log^{2}(T)T^{7/6}2^{-(n-2)}leqn^{2}\left(4\log^{2}(T)T^{7/6}\right)^{3/13}~{}\leq~{}4n^{2}\log(T)T^{7/26},$
and therefore
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t})$ $~{}\leq~{}4C\|\mathbf{C}\|\|\mathbf{B}\|n^{3}k\sqrt{d_{\textrm{out}}}\log(T)T^{10/13}.$
Observing that both $n$ and $k$ are bounded polylogarithmically in $T$ and dividing both sides by $T$ we get the stated result. ${1.5ex}{1.5ex}$ 0.0pt1.5ex1.5ex
Remark on the loss function.
The reader may notice we use the $\ell_{1}$, or absolute loss, rather than Euclidean or other loss functions. All norms are equivalent up to the (output) dimension, and thus learning to predict as well as the best linear dynamical system in hindsight is meaningful in any norm. However, we make this technical choice since it greatly simplifies the regret bounds, and in particular the bound on the gradient norms, which is technically involved. We conjecture that sublinear regret bounds are attainable in other norms as well, and leave it for future work.
We prove Theorem C.2 on an (equivalent) algorithm, where we rescale the parameter $\mathbf{M}_{j}$ by $\sqrt{T}$ and the input $\langle\phi_{j},\mathbf{u}_{(t-n-i):1}\rangle$ by $1/\sqrt{T}$. We account for this rescaling by increasing the size of the domain for $\mathbf{M}$ by $\sqrt{T}$. The proof of Theorem C.2 proceeds in two parts. The first is to show that any linear dynamical signal is well approximated by a predictor of the form in Algorithm 4 $\hat{\mathbf{y}}(\mathbf{Q},\mathbf{M})$ where $(\mathbf{Q},\mathbf{M})\in\mathcal{K}$.
Lemma C.3 (Approximation Lemma).
Suppose $\mathbf{y}_{1:T}$ evolves as a linear dynamical system characterized by $(\mathbf{A},\mathbf{B},\mathbf{C})$ as in Eq. 3. Consider domain
[…]
If $R_{Q}~{}\geq~{}\|\mathbf{C}\|\|\mathbf{B}\|\|\mathbf{c}\|_{1}$, $R_{M}~{}\geq~{}C\log(T)^{1/6}T^{7/6}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\|\mathbf{C}\mathbf{B}\|$ and $k~{}\geq~{}C\log^{2}\left(T\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\|\mathbf{C}\mathbf{B}\|/(d_{\textrm{out}}\epsilon)\right)$ (where $C>0$ is a universal constant) then there exists $(\hat{\mathbf{Q}}_{1},\dots,\hat{\mathbf{Q}}_{n},\hat{\mathbf{M}}_{1},\dots,\hat{\mathbf{M}}_{k})\in\mathcal{K}$ such that for prediction (as in Algorithm 4)
$\hat{\mathbf{y}}_{t}(\hat{\mathbf{Q}},\hat{\mathbf{M}})=-\sum_{i=1}^{n}c_{i}\mathbf{y}_{t-i}+\sum_{j=0}^{n}\hat{\mathbf{Q}}^{t}_{j}\mathbf{u}_{t-j}+\frac{1}{\sqrt{T}}\sum_{j=1}^{k}\hat{\mathbf{M}}_{j}^{t}\phi_{j}^{\top}\tilde{\mathbf{u}}_{t-n-1:1},$
we have
$\|\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\|_{1}~{}\leq~{}\epsilon.$
Lemma C.3 is the heart of the result and requires several pages of proofs and novel ideas to establish the result. We therefore dedicate Section C.1 to its proof.
The next result provides the regret of Online Gradient Descent when compared to the best $(\mathbf{Q}^{*},\mathbf{M}^{*})\in\mathcal{K}$.
Lemma C.4 (Online Gradient Descent).
Recall the domain $\mathcal{K}$ in Algorithm 4 be
[…]
Given coefficients $\mathbf{c}_{1:n}$, let $p_{n}^{\mathbf{c}}(x)=x^{n}+\mathbf{c}_{1}x^{n-1}+\dots+\mathbf{c}_{n}$ and let $r$ denote the maximum absolute value of the coefficients. The iterates $\hat{y}_{t}^{{\mathcal{A}}}$ output by Algorithm 4 satisfy
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})-\min_{(\mathbf{Q}^{*},\mathbf{M}^{*})\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{*},\mathbf{M}^{*})~{}\leq~{}\frac{3}{2}(nR_{Q}+kR_{M})(n+k)\sqrt{d_{\textrm{out}}}\sqrt{T}.$
Proof of Lemma C.4.
Let $G=\max_{t\in[T]}\|\nabla\mkern-2.5mu_{\mathbf{Q},\mathbf{M}}\ell_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})\|$ and let
$D=\max_{(\mathbf{Q}^{1},\mathbf{M}^{1}),(\mathbf{Q}^{2},\mathbf{M}^{2})\in\mathcal{K}}\|(\mathbf{Q}^{1},\mathbf{M}^{1})-(\mathbf{Q}^{2},\mathbf{M}^{2})\|.$
By Theorem 3.1 from [22],
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})-\min_{(\mathbf{Q}^{*},\mathbf{M}^{*})\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{*},\mathbf{M}^{*})~{}\leq~{}\frac{3}{2}GD\sqrt{T}.$
Therefore it remains to bound $G$ and $D$. By definition of $\mathcal{K}$ we have
$D~{}\leq~{}nR_{Q}+kR_{M}.$
For $G$ we compute the subgradient at any $\mathbf{Q}_{i}$ and $\mathbf{M}_{i}$. Note that, in general, $\nabla\mkern-2.5mu_{\mathbf{A}}\|\mathbf{A}\mathbf{x}-\mathbf{b}\|_{1}=\mathrm{sign}(\mathbf{A}\mathbf{x}-\mathbf{b})\mathbf{x}^{\top}$. Since $\|\mathbf{x}\mathbf{y}^{\top}\|_{F}~{}\leq~{}\|\mathbf{x}\|_{2}\|\mathbf{y}\|_{2}$ we have
$\|\nabla\mkern-2.5mu_{\mathbf{A}}f(\mathbf{A})\|_{F}~{}\leq~{}\sqrt{d}\|\mathbf{x}\|_{2},$
where $d$ is the dimension of $\mathbf{b}$. Using this and the assumption that for any $t\in[T]$, $\|\mathbf{u}_{t}\|_{2}~{}\leq~{}1$, we have
$\|\nabla\mkern-2.5mu_{\mathbf{M}_{j}}\ell_{t}(\mathbf{Q},\mathbf{M})\|$ $~{}\leq~{}\sqrt{d_{\textrm{out}}}\|\phi_{j}^{\top}\tilde{\mathbf{u}}_{t-n-1:1}T^{-1/2}\|$
$~{}\leq~{}\sqrt{d_{\textrm{out}}}\|\phi_{j}\|_{1}\|\tilde{\mathbf{u}}_{t-n-1:1}\|_{\infty}T^{-1/2}$
$~{}\leq~{}\sqrt{d_{\textrm{out}}}.$ ($\|\tilde{\mathbf{u}}_{t-n-1:1}\|_{\infty}~{}\leq~{}1$ and $\|\phi_{j}\|_{1}~{}\leq~{}\sqrt{T}$ since $\|\phi_{j}\|_{2}~{}\leq~{}1$ )
Next,
$\|\nabla\mkern-2.5mu_{\mathbf{Q}_{i}}\ell_{t}(\mathbf{Q})\|_{F}$ $~{}\leq~{}\sqrt{d_{\textrm{out}}}\|\mathbf{u}_{t-i}\|_{2}~{}\leq~{}\sqrt{d_{\textrm{out}}}.$
Therefore,
$G=\max_{t\in[T]}\|\nabla\mkern-2.5mu_{(\mathbf{Q},\mathbf{M})}\ell_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})\|_{F}~{}\leq~{}(n+k)\sqrt{d_{\textrm{out}}}.$
Therefore, we have
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})-\min_{(\mathbf{Q}^{*},\mathbf{M}^{*})\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{*},\mathbf{M}^{*})~{}\leq~{}\frac{3}{2}(nR_{Q}+kR_{M})(n+k)\sqrt{d_{\textrm{out}}}\sqrt{T}.$
${1.5ex}{1.5ex}$ 0.0pt1.5ex1.5ex
Combining Lemma C.3 and Lemma C.4 proves Theorem C.2.
Proof of Theorem C.2.
By Lemma C.4 the iterates from Algorithm 4 $(\mathbf{Q}^{1},\mathbf{M}^{1})$, …, $(\mathbf{Q}^{T},\mathbf{M}^{T})$ satisfy
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})-\min_{(\mathbf{Q}^{*},\mathbf{M}^{*})\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{*},\mathbf{M}^{*})~{}\leq~{}\frac{3}{2}(nR_{Q}+kR_{M})(n+k)\sqrt{d_{\textrm{out}}}\sqrt{T}.$
Next, if $\mathbf{y}_{1:T}$ comes from a linear dynamical system, by Lemma C.3, if $R_{Q}~{}\geq~{}\|\mathbf{C}\|\|\mathbf{B}\|\|\mathbf{c}\|_{1}$, $R_{M}~{}\geq~{}C\log(T)^{1/6}T^{7/6}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\|\mathbf{C}\mathbf{B}\|$ and $k~{}\geq~{}C\log^{2}\left(T\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\|\mathbf{C}\mathbf{B}\|/(d_{\textrm{out}}\epsilon)\right)$ (where $C>0$ is a universal constant) then there exists $(\hat{\mathbf{Q}},\hat{\mathbf{M}})\in\mathcal{K}$ such that
$\sum_{t=1}^{T}\ell_{t}(\hat{\mathbf{Q}},\hat{\mathbf{M}})~{}\leq~{}\epsilon T.$
Since $(\hat{\mathbf{Q}},\hat{\mathbf{M}})\in\mathcal{K}$ we must also have that
$\min_{(\mathbf{Q}^{*},\mathbf{M}^{*})\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{*},\mathbf{M}^{*})~{}\leq~{}\epsilon T.$
Therefore,
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})$ $~{}\leq~{}\min_{(\mathbf{Q}^{*},\mathbf{M}^{*})\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{*},\mathbf{M}^{*})+\frac{3}{2}(nR_{Q}+kR_{M})(n+k)\sqrt{d_{\textrm{out}}}\sqrt{T}$
$~{}\leq~{}\epsilon T+\frac{3}{2}(nR_{Q}+kR_{M})(n+k)\sqrt{d_{\textrm{out}}}\sqrt{T}.$
Plugging in the values for $R_{Q}$ and $R_{M}$ and for $k$ when $\epsilon=1/T$ we conclude
$\sum_{t=1}^{T}\ell_{t}(\mathbf{Q}^{t},\mathbf{M}^{t})$ $~{}\leq~{}\epsilon T+\frac{3}{2}\|\mathbf{C}\|\|\mathbf{B}\|\bigg{(}n\|\mathbf{c}\|_{1}$
$\qquad+C\log^{2}\left(T\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}(\alpha)|\|\mathbf{C}\|\|\mathbf{B}\|\right)T^{7/6}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}(\alpha)}|\bigg{)}(n+k)\sqrt{d_{\textrm{out}}}\sqrt{T}$
$~{}\leq~{}1+C\|\mathbf{C}\|\|\mathbf{B}\|\left(n\|\mathbf{c}\|_{1}+\log^{2}(T)T^{7/6}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\right)(n+k)\sqrt{d_{\textrm{out}}}\sqrt{T}.$
${1.5ex}{1.5ex}$ 0.0pt1.5ex1.5ex
C.1 Proving Approximation Lemma C.3
The difficult result to prove is Lemma C.3. As discussed in Section 4, any signal from a linear dynamical system can be broken down into two parts: a linear autoregressive term and a spectral filtering term (see Eq (7)). The “spectral filtering term” is $\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s}$. We call it as such because its structure is amenable to spectral filtering. Indeed, we construct a set of spectral filters that satisfy the following.
Lemma C.5.
Let $\mu(\alpha)$ be as defined in Eq (1) and $\mathbf{Z}$ be as defined in Eq (2) and let $\phi_{1},\dots,\phi_{T-n}$ be the eigenvectors of $\mathbf{Z}_{T-n}$. There is a universal constant $C>0$ such that
$\max_{\alpha\in\mathbb{C}_{\beta}}|\mu(\alpha)^{\top}\phi_{j}|~{}\leq~{}C\log(T)^{1/6}T^{2/3}c^{-j/6\log T}.$
The proof of Lemma C.5 is inspired by previous spectral filtering literature but requires several key and nontrivial adaptations to incorporate the polynomial $p_{n}^{\mathbf{c}}(\cdot)$, to argue that the new spectral filtering matrix $\mathbf{Z}$ defined over a complex domain exhibits exponential eigenvalue decay, and to generally extend the analysis to the case where the spectrum of $\mathbf{A}$ may lie in the complex plane. We prove both lemmas in the subsequent subsections. With Lemma C.5 in hand, we may prove Lemma C.3.
Proof of Lemma C.3.
Suppose $\mathbf{y}_{1:T}$ evolves as a linear dynamical system characterized by $(\mathbf{A},\mathbf{B},\mathbf{C})$. Then given inputs $\mathbf{u}_{1:t}$,
$\mathbf{y}_{t}=\sum_{s=1}^{t}\mathbf{C}\mathbf{A}^{t-s}\mathbf{B}\mathbf{u}_{s}.$
With some linear algebra we get that convolving $\mathbf{y}_{t:t-n}$ with coefficients $\mathbf{c}_{0:n}$ results in
$\sum_{i=0}^{n}\mathbf{c}_{i}\mathbf{y}_{t-i}=\sum_{s=0}^{n-1}\sum_{i=0}^{s}\mathbf{c}_{i}\mathbf{C}\mathbf{A}^{s-i}\mathbf{B}\mathbf{u}_{t-s}+\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}^{\mathbf{c}}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s},$
or equivalently (since $\mathbf{c}_{0}=1$ due to the assertion in Algorithm 3),
$\mathbf{y}_{t}=-\sum_{i=1}^{n}\mathbf{c}_{i}\mathbf{y}_{t-i}+\sum_{s=0}^{n-1}\sum_{i=0}^{s}\mathbf{c}_{i}\mathbf{C}\mathbf{A}^{s-i}\mathbf{B}\mathbf{u}_{t-s}+\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}^{\mathbf{c}}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s}.$ (8)
Set $\hat{\mathbf{Q}}_{s}=\sum_{i=0}^{s}\mathbf{c}_{i}\mathbf{C}\mathbf{A}^{s-i}\mathbf{B}$ and set $\hat{\mathbf{Q}}=\left(\hat{\mathbf{Q}}_{0},\dots,\hat{\mathbf{Q}}_{n-1}\right)$. Then
$\|\hat{\mathbf{Q}}_{i}\|~{}\leq~{}\sum_{j=0}^{i}|\mathbf{c}_{j}|\|\mathbf{C}\mathbf{A}^{i-j}\mathbf{B}\|~{}\leq~{}\|\mathbf{C}\|\|\mathbf{B}\|\sum_{j=0}^{i}|\mathbf{c}_{j}|~{}\leq~{}\|\mathbf{C}\|\|\mathbf{B}\|\|\mathbf{c}\|_{1}.$
Next we turn our attention to the spectral filtering parameters. Recall our definition of $\mu_{p_{n}}(\alpha)$,
$\mu(\alpha)\stackrel{{\scriptstyle\text{def}}}{{=}}\begin{bmatrix}1&\alpha&\dots&\alpha^{t-n-1}\end{bmatrix}^{\top}$
and
$\mathbf{u}_{t:1}\stackrel{{\scriptstyle\text{def}}}{{=}}\begin{bmatrix}&\mathbf{u}_{t}&\mathbf{u}_{t-1}&\dots&\mathbf{u}_{1}\end{bmatrix}^{\top}.$ (9)
Let $\mathbf{A}$ be diagonalized by some $\mathbf{P}$ so that
$\mathbf{A}=\mathbf{P}\mathbf{D}\mathbf{P}^{-1},$
where $\mathbf{D}$ is the diagonalization of $\mathbf{A}$, note that we can assume this w.l.o.g. since the set of diagonalizable matrices over the complex numbers is dense and therefore if $\mathbf{A}$ is not diagonalizable we may apply an arbitrarily small perturbation to it. Recall the notation from Algorirthm 4 that $\tilde{\mathbf{u}}_{t:1}$ is $\mathbf{u}_{t:1}$ padded so that it has shape $(T-n-1)\times d_{\textrm{in}}$. Then,
$\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}^{\mathbf{c}}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s}$ $=\sum_{s=0}^{t-n-1}\mathbf{C}\mathbf{P}p_{n}^{\mathbf{c}}(\mathbf{D})\mathbf{D}^{s}\mathbf{P}^{-1}\mathbf{B}\mathbf{u}_{t-n-s}$
$=\sum_{s=0}^{t-n-1}\mathbf{C}\mathbf{P}\left(\sum_{m=1}^{d_{h}}p_{n}^{\mathbf{c}}(\alpha_{m})\alpha_{m}^{s}\mathbf{e}_{m}\mathbf{e}_{m}^{\top}\right)\mathbf{P}^{-1}\mathbf{B}\mathbf{u}_{t-n-s}$
$=\sum_{m=1}^{d_{h}}\sum_{s=0}^{t-n-1}\mathbf{C}\mathbf{P}\mathbf{e}_{m}\mathbf{e}_{m}^{\top}\mathbf{P}^{-1}\mathbf{B}p_{n}^{\mathbf{c}}(\alpha_{m})\alpha_{m}^{s}\mathbf{u}_{t-n-s}$
$=\sum_{m=1}^{d_{h}}\mathbf{C}\mathbf{P}\mathbf{e}_{m}\mathbf{e}_{m}^{\top}\mathbf{P}^{-1}\mathbf{B}\mu_{p_{n}}(\alpha_{m})^{\top}\tilde{\mathbf{u}}_{t-n-1:1}$
$=\sum_{m=1}^{d_{h}}\mathbf{C}\mathbf{P}\mathbf{e}_{m}\mathbf{e}_{m}^{\top}\mathbf{P}^{-1}\mathbf{B}\mu_{p_{n}}(\alpha_{m})^{\top}\left(\sum_{j=1}^{T-n}\phi_{j}\phi_{j}^{\top}\right)\tilde{\mathbf{u}}_{t-n-1:1}$ (Orthonormality of the filters)
$=\sum_{j=1}^{T-n}\left(\sum_{m=1}^{d_{h}}\mathbf{C}\mathbf{P}\mathbf{e}_{m}\mathbf{e}_{m}^{\top}\mathbf{P}^{-1}\mathbf{B}\mu_{p_{n}}(\alpha_{m})^{\top}\phi_{j}\right)\phi_{j}^{\top}\tilde{\mathbf{u}}_{t-n-1:1}.$
Therefore defining
$\hat{\mathbf{M}}_{j}\stackrel{{\scriptstyle\text{def}}}{{=}}T^{1/2}\sum_{m=1}^{d_{h}}\mathbf{C}\mathbf{P}\mathbf{e}_{m}\mathbf{e}_{m}^{\top}\mathbf{P}^{-1}\mathbf{B}\mu_{p_{n}}(\alpha_{m})^{\top}\phi_{j},$
we have
$\sum_{s=0}^{t-n-1}\mathbf{C}p_{n}^{\mathbf{c}}(\mathbf{A})\mathbf{A}^{s}\mathbf{B}\mathbf{u}_{t-n-s}$ $=\sum_{j=1}^{T-n}\hat{\mathbf{M}}_{j}\frac{\phi_{j}^{\top}\tilde{\mathbf{u}}_{t-n-1:1}}{\sqrt{T}}.$
Next we bound the norm of $\hat{\mathbf{M}}_{j}$. Recalling that $\mu_{p_{n}^{\mathbf{c}}}(\alpha)=p_{n}^{\mathbf{c}}(\alpha)\mu(\alpha)$,
$\|\hat{\mathbf{M}}_{j}\|$ $=\|T^{1/2}\sum_{m=1}^{d_{h}}\mathbf{C}\mathbf{P}\mathbf{e}_{m}\mathbf{e}_{m}^{\top}\mathbf{P}^{-1}\mathbf{B}p_{n}^{\mathbf{c}}(\alpha_{m})\mu(\alpha_{m})^{\top}\phi_{j}\|$
$~{}\leq~{}T^{1/2}\max_{m\in[d_{h}]}|p_{n}^{\mathbf{c}}(\alpha_{m})|\cdot\|\mathbf{C}\mathbf{B}\|\cdot\max_{m\in[d_{h}]}|\mu(\alpha_{m})^{\top}\phi_{j}|$
$~{}\leq~{}T^{1/2}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\cdot\|\mathbf{C}\mathbf{B}\|\cdot\max_{\alpha\in\mathbb{C}_{\beta}}|\mu(\alpha)^{\top}\phi_{j}|$ (The spectrum of $\mathbf{A}$ lies in $\mathbb{C}_{\beta}$ by assumption.)
By Lemma C.5, there are universal constants $C>0$ and $c>1$ such that
$\max_{\alpha\in\mathbb{C}_{\beta}}|\mu(\alpha)^{\top}\phi_{j}|~{}\leq~{}C\log(T)^{1/6}T^{2/3}c^{-j/6\log T}.$
Therefore,
$\|\hat{\mathbf{M}}_{j}\|~{}\leq~{}C\log(T)^{1/6}T^{7/6}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\cdot\|\mathbf{C}\mathbf{B}\|\cdot c^{-j/6\log T}.$
Therefore $(\hat{\mathbf{Q}}_{0},\dots,\hat{\mathbf{Q}}_{n-1},\hat{\mathbf{M}}_{1},\dots,\hat{\mathbf{M}}_{k})\in\mathcal{K}$ as long as $R_{Q}~{}\geq~{}\|\mathbf{C}\|\|\mathbf{B}\|\|\mathbf{c}\|_{1}$ and $R_{M}\geq C\log(T)^{1/6}T^{7/6}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\|\mathbf{C}\mathbf{B}\|$. Therefore, recalling Eq. 8, we have shown,
$\mathbf{y}_{t}=-\sum_{i=1}^{n}\mathbf{c}_{i}\mathbf{y}_{t-i}+\sum_{s=0}^{n-1}\sum_{i=0}^{s}\hat{\mathbf{Q}}_{s}\mathbf{u}_{t-s}+\sum_{j=1}^{T-n}\hat{\mathbf{M}}_{j}\frac{\phi_{j}^{\top}\tilde{\mathbf{u}}_{t-n-1:1}}{\sqrt{T}}.$
Observe that $\hat{\mathbf{y}}(\hat{\mathbf{Q}},\hat{\mathbf{M}})$ is nearly identical to the above, but truncates the terms involving $\hat{\mathbf{M}}_{j}$ after $j>k$, where $k$ is a fixed parameter input to Algorithm 4. Therefore,
$\hat{\mathbf{y}}_{t}(\hat{\mathbf{Q}},\hat{\mathbf{M}})-\mathbf{y}_{t}=\sum_{j=k+1}^{T-n}\hat{\mathbf{M}}_{j}\frac{\phi_{j}^{\top}\tilde{\mathbf{u}}_{t-n-1:1}}{\sqrt{T}}.$
Therefore, to bound the error we have
$\|\hat{\mathbf{y}}_{t}(\hat{\mathbf{Q}},\hat{\mathbf{M}})-\mathbf{y}_{t}\|_{2}$ $=\|\sum_{j=k+1}^{T-n}\mathbf{M}_{j}\phi_{j}^{\top}\mathbf{u}_{(t-n-1):1}T^{-1/2}\|_{2}$
$~{}\leq~{}\sum_{j=k+1}^{T-n}\|\hat{\mathbf{M}}_{j}\|\|\phi_{j}^{\top}\mathbf{u}_{(t-n-1):1}T^{-1/2}\|_{2}$
$~{}\leq~{}\sum_{j=k+1}^{T-n}\|\hat{\mathbf{M}}_{j}\|\|\phi_{j}^{\top}\|_{1}\|\mathbf{u}_{(t-n-1):1}\|_{\infty}T^{-1/2}$
$~{}\leq~{}\sum_{j=k+1}^{T-n}\|\hat{\mathbf{M}}_{j}\|$ ($\|\phi_{j}\|_{2}=1$ and therefore $\|\phi_{j}\|_{1}~{}\leq~{}\sqrt{T}$ )
$~{}\leq~{}C\log(T)^{1/6}T^{1+(7/6)}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\|\mathbf{C}\mathbf{B}\|c^{-k/6\log T}$
$~{}\leq~{}\epsilon/\sqrt{d_{\textrm{out}}}.$ (Whenever $k~{}\geq~{}6\log(T)\log\left(C\log(T)^{1/6}T^{2}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\|\mathbf{C}\mathbf{B}\|/\sqrt{d_{\textrm{out}}}\epsilon\right)/\log(c)$.)
Therefore as long as $C>0$ is such that
$k\stackrel{{\scriptstyle\text{def}}}{{=}}C\log^{2}\left(T\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\|\mathbf{C}\mathbf{B}\|/(d_{\textrm{out}}\epsilon)\right)~{}\geq~{}6\log(T)\log\left(C\log(T)^{1/6}T^{2}\max_{\alpha\in\mathbb{C}_{\beta}}|p_{n}^{\mathbf{c}}(\alpha)|\|\mathbf{C}\mathbf{B}\|/\sqrt{d_{\textrm{out}}}\epsilon\right)/\log(c),$
the statement above holds. Finally,
$\|\hat{\mathbf{y}}_{t}(\hat{\mathbf{Q}},\hat{\mathbf{M}})-\mathbf{y}_{t}\|_{1}~{}\leq~{}\sqrt{d_{\textrm{out}}}\|\hat{\mathbf{y}}_{t}(\hat{\mathbf{Q}},\hat{\mathbf{M}})-\mathbf{y}_{t}\|_{2}~{}\leq~{}\epsilon.$
${1.5ex}{1.5ex}$ 0.0pt1.5ex1.5ex
C.1.1 Proof of the Spectral Filtering Property Lemma C.5
In order to prove Lemma C.5 we require two further helper lemmas. The first is Lemma C.6, which roughly argues that the Lipschitz constant of a function $f:\mathbb{C}_{\beta}\to{\mathcal{R}}$ can be bounded by a polynomial of the expectation of the function on $\mathbb{C}_{\beta}$. The second is Theorem C.7, which argues that the matrix $\mathbf{Z}$ from Eq. 2 that we use to derive the new spectral filters exhibits exponentially fast decay in its spectrum. This is a nontrivial and novel proof that does not follow from previous spectral filtering literature and, as such, we have chosen to use the word “Theorem” rather than “Lemma” to refer to it.
Lemma C.6.
Let $L>0$, $g_{\max}>0$ and $0\leq\beta\leq 1$. Define
$\mathbb{C}_{\beta}=\bigl{\{}\,z\in\mathbb{C}:\;|z|\leq 1,\;|\arg(z)|\leq\beta\bigr{\}},$
and let $\mathcal{F}$ be the set of non-negative, $L$-Lipschitz functions $f:\mathbb{C}_{\beta}\!\to\!\mathbb{R}$ such that $\max_{z\in\mathbb{C}_{\beta}}f(z)=g_{\max}$. Then
[…]
Proof of Lemma C.6.
1. The extremal function. Fix $f\in\mathcal{F}$ and choose $z_{\star}$ with $f(z_{\star})=g_{\max}$. Among admissible functions we consider only those that decrease as fast as the Lipschitz constraint allows in every radial direction, replacing $f$ by the extremal cone. Set
$h(z):=[\,g_{\max}-L|z-z_{\star}|\,]_{+},\qquad r:=\frac{g_{\max}}{L}.$
Because $0\leq h\leq f$ on $\mathbb{C}_{\beta}$, it suffices to lower bound $\int_{\mathbb{C}_{\beta}}h$.
2. Geometry around $z_{\star}$. It can be seen by Euclidean geometry, the point $z_{\star}=0$ minimize the volume of the intersection with $C_{\beta}$, up to a factor of 44{4}4, to make it the extremal function. It is depicted in Figure 1. Notice that the area of the intersection is colored in blue, and it is particularly easy to integrate over.
Figure 1: The origin minimizes the intersection area up to factor $\frac{1}{4}$.
3. Integrating $h$.
[…]
Since $h\in\mathcal{F}$, and is the extremal function, we obtain the Lemma statement. ${1.5ex}{1.5ex}$ 0.0pt1.5ex1.5ex
Theorem C.7.
Given horizon $T$ let
$\mu(\alpha)\stackrel{{\scriptstyle\text{def}}}{{=}}\begin{bmatrix}1&\alpha&\dots&\alpha^{T-1}\end{bmatrix}^{\top},$
and
[…]
where $\overline{\alpha}\in\mathbb{C}$ denotes the complex conjugate and $\mathbb{C}_{\beta}=\left\{z\in\mathbb{C}\textrm{ s.t. }|z|~{}\leq~{}1\textrm{ and }|\mathrm{arg}(z)|~{}\leq~{}\theta\right\}$. There exist constants $C>0$ and $c>1$, independent of $T$, such that if $\sigma_{j}(\mathbf{Z})$ denotes the $j$-th largest singular value of $\mathbf{Z}$,
$\sigma_{j}~{}\leq~{}C\beta\cdot c^{-j/\log T}\cdot\log T.$
This proof is nontrivial and does not follow from previous results on Spectral Filtering due to the fact that it is no longer a Hankel matrix. It also does not immediately follow from general work on spectral decay, for instance [11], since it does not have an obvious displacement structure. We prove its spectral decay borrowing heavily from techniques from [11], but we must introduce a novel Gaussian quadrature argument that applies to the complex plane.
Proof of Theorem C.7.
Observe that we can write $\mathbf{Z}_{jk}$ as
[…]
or, equivalently
[…]
$=\int_{r\in[0,1]}\int_{\theta\in\mathbb{C}_{\beta}}r^{j+k+1}\cos(|j-k|\theta)drd\theta+i\cdot\mathrm{sign}(j-k)\int_{r\in[0,1]}\int_{\theta\in\mathbb{C}_{\beta}}\sin(|j-k|)\theta)drd\theta.$
We will use Gaussian quadrature in what follows. We have two non-standard features that we must manage. First, it is commonly known that Gaussian quadrature with $n$ nodes is exact when integrating a polynomial of degree $12n-$ 1 or less. However, we need to ensure exactness when integrating a trigonometric function. The second non-standard feature is that we are integrating a $2$-dimensional function, ie our domain is $2$-dimensional specified by $(r,\theta)$, while standard Gaussian quadrature applies only to $1$-dimensional functions.
For the first non-standard feature, [27] shows that Guassian quadrature is still exact even when integrating trigonometric functions. Specifically, they show that if the domain of integration is some compact set $\mathcal{C}$, then there there exist $n+1$ weights $w_{1}^{\mathrm{trig}},\dots,w_{n}^{\mathrm{trig}}$ and nodes $\theta_{1},\dots,\theta_{n}$ such that for any $f(\theta)$ which can be written in the form
$f(\theta)=\sum_{\ell=1}^{n}a_{\ell}\sin(\ell\theta)+\sum_{\ell=0}^{n}b_{\ell}\cos(\ell\theta),$ (10)
we have
[…]
Let $f_{1}^{(j,k)}(r,\theta)=r^{j+k+1}\cos(|j-k|\theta)$ and $f_{2}^{(j,k)}(r,\theta)=r^{j+k+1}\sin(|j-k|\theta)$. Using the above, observe that both $f_{1}^{(j,k)}$ and $f_{2}^{(j,k)}$ can be written as polynomials in $r$ and $\theta$ with degree at most $2|j-k|~{}\leq~{}2T$.
For the second nonstandard feature, i.e. that we are integrating over a $2$-dimensional function, we can use Tchakaloff’s Theorem, which is an extension of quadrature rules to general Borel measures on $\mathbb{R}^{d}$. Given any positive Borel measure on ${\mathcal{R}}^{d}$, $\mu$, which has compact support, Tchakaloff’s Theorem guarantees the existence of $n+1$ weights $w_{1},\dots,w_{n+1}$ and nodes $v_{1},\dots,v_{n+1}$ such that for any polynomial $P(x)$ on ${\mathcal{R}}^{d}$ with degree at most $n$,
$\int_{{\mathcal{R}}^{d}}P(x)\mu(dx)=\sum_{i=1}^{n+1}w_{i}P(v_{i}).$ (12)
Since our measure is compactly supported (we have $drd\theta$ integrated over a compact domain) we have that there exist weights $w_{1}^{2},\dots,w_{2T+1}^{2}$ and nodes $(r_{1},\theta_{1}),\dots,(r_{2T+1},\theta_{2T})$ such that for any $j,k~{}\leq~{}T$ both $f_{1}^{(j,k)}(r,\theta)$ and $f_{2}^{(j,k)}(r\theta)$ satisfy
$\int_{r\in[0,1]}\int_{\theta\in\mathbb{C}_{\beta}}f_{i}^{(j,k)}(r,\theta)=\sum_{\ell=1}^{2T}w_{\ell}f_{1}^{(j,k)}(r_{\ell},\theta_{\ell}).$
Therefore
$\mathbf{Z}_{jk}$ $=\sum_{\ell=1}^{2T+1}\left(w_{\ell}^{2}r_{\ell}^{j+k+1}\cos(|j-k|\theta_{\ell})\right)+i\text{ }\mathrm{sign}(j-k)\left(\sum_{\ell=1}^{n+1}w_{\ell}^{2}r_{\ell}^{j+k+1}\sin(|j-k|\theta_{\ell})\right)$
$=$ $\sum_{\ell=1}^{2T+1}w_{\ell}^{2}r_{\ell}^{j+k+1}\left(\cos((j-k)\theta_{\ell})+i\sin((j-k)\theta_{\ell})\right)$
$=$ $\sum_{\ell=1}^{2T+1}w_{\ell}^{2}r_{\ell}^{j+k+1}e^{i(j-k)\theta_{\ell}}$
$=$ $\sum_{\ell=1}^{2T+1}w_{\ell}^{2}\left(r_{\ell}e^{i\theta_{\ell}}\right)^{j}\left(r_{\ell}e^{-i\theta_{\ell}}\right)^{k}$
$=$ $\sum_{\ell=1}^{2T+1}w_{\ell}^{2}z_{\ell}^{j}\bar{z}_{\ell}^{k},$
where $z_{\ell}\stackrel{{\scriptstyle\text{def}}}{{=}}r_{\ell}e^{i\theta_{\ell}}$. We now proceed by using the same proof technique from [11]. Indeed, we have shown that $\mathbf{Z}$ has a so-called “Fiedler factorization”,
$\mathbf{Z}=\mathbf{K}\mathbf{K}^{*},$ (13)
where
$\mathbf{w}$ $=\begin{bmatrix}w_{1}&\dots&w_{2T+1}\end{bmatrix},$
$\mathbf{z}$ $=\begin{bmatrix}z_{1}&\dots&z_{2T+1}\end{bmatrix},$
$\mathbf{K}$ $=\begin{bmatrix}\mathbf{w}&\mathrm{diag}(\mathbf{z})\mathbf{w}&\dots&\left(\mathrm{diag}(\mathbf{z})\right)^{2T-1}\mathbf{w}\end{bmatrix}.$
Indeed,
$\left(\mathbf{K}\mathbf{K}^{*}\right)_{jk}$ $=\sum_{\ell=1}^{2T-1}\left(\mathbf{K}\right)_{j\ell}\mathbf{K}^{*}_{\ell k}$
$=\sum_{\ell=1}^{2T-1}\left(\mbox{diag}(\mathbf{z})^{j}\mathbf{w}\right)_{\ell}\left(\mbox{diag}(\mathbf{z}^{*})^{k}\mathbf{w}\right)_{\ell}$
$=\sum_{\ell=1}^{2T-1}\left(\mbox{diag}(\mathbf{z})^{j}\mathbf{w}\right)_{\ell}\left(\mbox{diag}(\mathbf{z}^{*})^{k}\mathbf{w}\right)_{\ell}$
$=\sum_{\ell=1}^{2T-1}z^{j}_{\ell}w_{\ell}\overline{z}_{\ell}^{k}w_{\ell}$
$=\mathbf{Z}_{jk}.$
Recall that an $m\times n$ matrix $\mathbf{K}_{m,n}$ with $m~{}\geq~{}n$ is said to be a Krylov matrix with Hermitian argument if there exists a Hermitian matrix $\mathbf{A}\in\mathbb{C}^{m\times m}$ and a vector $w\in\mathbb{C}^{m}$ such that
$\mathbf{K}_{m,n}=\begin{bmatrix}w&\mathbf{A}w&\dots&\mathbf{A}^{n-1}w\end{bmatrix}.$
We have the following lemma which comes from Corollary 5.2 of [11].
Lemma C.8.
[Corollary 5.2 of [11]] Let $\mathbf{K}_{m,n}$ be an $m\times n$ Krylov matrix. The singular values of $\mathbf{K}_{m,n}$ can be bounded as follows:
$\sigma_{j+2k}\left(\mathbf{K}_{m,n}\right)~{}\leq~{}4\left[\exp\left(\frac{\pi^{2}}{4\log(8\lfloor n/2\rfloor/\pi)}\right)\right]^{-k+1}\sigma_{j}\left(\mathbf{K}_{m,n}\right),\qquad 1~{}\leq~{}j+2k~{}\leq~{}n.$
Recall Eq (13) that $\mathbf{Z}=\mathbf{K}\mathbf{K}^{*}$. Thus the singular values of $\mathbf{Z}$ are directly related to the singular values of $\mathbf{K}$: $\sigma_{j}(\mathbf{Z})=\sigma_{j}(\mathbf{K})^{2}$. Since $\mathbf{K}$ is a $(2T+1)\times(2T-1)$ Krylov matrix, we conclude
$\sigma_{j+2k}(\mathbf{Z})$ $=\sigma_{j+2k}(\mathbf{K})^{2}$
$~{}\leq~{}16\left[\exp\left(\frac{\pi^{2}}{4\log(8\lfloor(2T+1)/2\rfloor/\pi)}\right)\right]^{-2k+2}\sigma_{j}\left(\mathbf{K}\right)^{2}$
$=16\left[\exp\left(\frac{\pi^{2}}{4\log(8T/\pi)}\right)\right]^{-2k+2}\sigma_{j}\left(\mathbf{Z}\right).$
Note that $\log(8T/\pi)~{}\leq~{}2\log(T)$ for $T>2$ and therefore if we set
$c\stackrel{{\scriptstyle\text{def}}}{{=}}\exp\left(\frac{\pi^{2}}{8}\right)\qquad\textrm{ and }\qquad C\stackrel{{\scriptstyle\text{def}}}{{=}}16\exp\left(\frac{\pi^{2}}{4}\right)^{2},$
we have
$16\left[\exp\left(\frac{\pi^{2}}{4\log(8T/\pi)}\right)\right]^{-2k+2}~{}\leq~{}C\cdot c^{-2k/\log T}.$
Given any index $m>2$, either $m$ is odd, $m=1+2k$, or $m$ is even, $m=2+2k$. Then
$\sigma_{m}=\sigma_{1+2k}~{}\leq~{}C\cdot c^{-2k/\log T}\cdot\sigma_{1},$
or
$\sigma_{m}=\sigma_{2+2k}~{}\leq~{}C\cdot c^{-2k/\log T}\cdot\sigma_{2}.$
Either way, we can bound $\max\left\{\sigma_{1},\sigma_{2}\right\}$ by $\mathrm{tr}(\mathbf{Z})$. Recall
[…]
Evaluating this integral for $j=k$ we get
$\mathbf{Z}_{jj}=\frac{4\beta}{2j+2}.$
Therefore,
$\mathrm{tr}(\mathbf{Z})$ $=4\beta\sum_{j=1}^{T}\frac{1}{2(j+1)}~{}\leq~{}2\beta\log(T).$
Plugging this in, we conclude that for any $j\in[T]$,
$\sigma_{j}~{}\leq~{}C\beta\cdot c^{-j/\log T}\cdot\log T.$
${1.5ex}{1.5ex}$ 0.0pt1.5ex1.5ex
With Lemma C.6 and Theorem C.7 in hand, we are ready to prove Lemma C.5.
Proof of Lemma C.5.
Let
$f_{j}(\alpha)\stackrel{{\scriptstyle\text{def}}}{{=}}|\phi_{j}^{\top}\mu(\alpha)|^{2},$
If $f_{j}(\alpha)$ is $L$-Lipschitz, letting $g_{\textrm{max}}=\max_{\alpha\in\mathbb{C}_{\beta}}f_{j}(\alpha)$ by Lemma C.6,
[…]
or equivalently,
[…]
Observe that
[…]
$=\sigma_{j}.$
Therefore conclude we have shown,
$\max_{\alpha\in\mathbb{C}_{\beta}}|\mu(\alpha)^{\top}\phi_{j}|=\max_{\alpha\in\mathbb{C}_{\beta}}\sqrt{f_{j}(\alpha)}~{}\leq~{}\left(\frac{24L^{2}}{\beta}\sigma_{j}\right)^{1/6}.$ (14)
The remainder of the proof consists of bounding the Lipschitz constant $L$ and bounding the eigenvalue $\sigma_{j}$. To bound the Lipschitz constant of $f_{j}$,
$L$ $~{}\leq~{}\max_{\alpha\in S}|f^{\prime}_{j}(\alpha)|$
$=\max_{\alpha\in S}2|\mathrm{Re}\left(\phi_{j}^{\top}\mu^{\prime}(\alpha)\cdot\phi_{j}^{\top}\mu(\alpha)\right)|$
$~{}\leq~{}\max_{\alpha\in S}\|\phi_{j}\|_{2}^{2}\cdot\|\mu(\alpha)\|_{2}\cdot\|\mu^{\prime}(\alpha)\|_{2}$
$=\max_{\alpha\in S}\|\mu(\alpha)\|_{2}\cdot\|\mu^{\prime}(\alpha)\|_{2}.$
We have
$\|\mu(\alpha)\|_{2}^{2}$ $=\left(\sum_{s=0}^{T-n-1}|\alpha|^{2s}\right)$
$~{}\leq~{}T.$
Next, to bound $\|\mu^{\prime}(\alpha)\|_{2}$ we observe,
$\left(\mu^{\prime}(\alpha)\right)_{j}$ $=\frac{d}{d\alpha}\alpha^{j}=j\alpha^{j-1}\mathbf{1}_{j>0}.$
Therefore, $|\left(\mu^{\prime}(\alpha)\right)_{j}|~{}\leq~{}T$ and $\|\mu^{\prime}(\alpha)\|_{2}^{2}~{}\leq~{}T^{3}$. Thus,
$L^{2}~{}\leq~{}T^{4}.$ (15)
By Theorem C.7, there are universal constants $C>0$ and $c>1$ such that
$\sigma_{j}~{}\leq~{}C\beta\cdot c^{-j/\log T}\cdot\log T.$
Plugging this in, we get for $C^{\prime}=24C$,
$\max_{\alpha\in\mathbb{C}_{\beta}}|\mu(\alpha)^{\top}\phi_{j}|~{}\leq~{}\left(24CT^{4}c^{-j/\log T}\log T\right)^{1/6}=C^{\prime}\log(T)^{1/6}T^{2/3}c^{-j/6\log T}.$
${1.5ex}{1.5ex}$ 0.0pt1.5ex1.5ex
Appendix D Chebyshev Polynomials Evaluated in the Complex Plane
In this section we let $T_{n}$ denote the $n$-th Chebyshev polynomial and let $M_{n}$ denote the monic form.
Proof of Lemma 4.1.
We use that $M_{n}(z)=T_{n}(z)/2^{n-1}$ and
$T_{n}(z)=\cos\left(n\arccos(z)\right).$ (16)
If $|\mathrm{Im}(z)|~{}\leq~{}1/64n^{2}$ then by Lemma D.2, $\arccos(z)~{}\leq~{}1/n$. Therefore $n\arccos(z)~{}\leq~{}1$ and so by Fact D.1,
$T_{n}(z)=\cos(n\arccos(z))~{}\leq~{}2$ (17)
Now we turn to the derivative $M_{n}^{\prime}(z)$. It’s a fact that
$M_{n}^{\prime}(z)=\frac{n}{2^{n-1}}U_{n-1}(z),$ (18)
where $U_{n-1}$ is the Chebyshev polynomial of the second kind. We next use the fact that
$U_{n-1}(z)=\begin{cases}2\sum_{\begin{subarray}{c}j~{}\geq~{}0\\ j\textrm{ even }\end{subarray}}^{n}T_{j}(z),&n\textrm{ even},\\ 2\sum_{\begin{subarray}{c}j~{}\geq~{}0\\ j\textrm{ odd }\end{subarray}}^{n}T_{j}(z),&n\textrm{ odd}.\end{cases}$
By Eq (17), $|T_{j}(z)|~{}\leq~{}2$ for any $j$ and therefore
$|U_{n-1}(z)|~{}\leq~{}n.$ (19)
Therefore,
$|M_{n}^{\prime}(z)|~{}\leq~{}\frac{n^{2}}{2^{n-1}}.$
${1.5ex}{1.5ex}$ 0.0pt1.5ex1.5ex
Fact D.1.
Let $z\in\mathbb{C}$. Then $|\cos(z)|~{}\leq~{}2$ whenever $|\mathrm{Im}|~{}\leq~{}1$.
Proof of Fact D.1.
$|\cos(x+iy)|$ $=\left(\cos^{2}x\cosh^{2}y+\sin^{2}x\sinh^{2}y\right)^{1/2}$ (Uses standard complex cosine identity.)
$=\left(\cos^{2}x+\cos^{2}x\left(\cosh^{2}y-1\right)+\sin^{2}x\sinh^{2}y\right)^{1/2}$
$=\left(\cos^{2}x+\cos^{2}x\sinh^{2}y+\sin^{2}x\sinh^{2}y\right)^{1/2}$ ($\cosh^{2}y-\sinh^{2}y=1$)
$=\left(\cos^{2}x+\sinh^{2}y\right)^{1/2}$ ($\cos^{2}x+\sin^{2}x=1$)
$~{}\leq~{}\left(1+\sinh^{2}y\right)^{1/2}$ ($\sinh^{2}y~{}\leq~{}2$ when $|y|~{}\leq~{}1$.)
$~{}\leq~{}2.$
${1.5ex}{1.5ex}$ 0.0pt1.5ex1.5ex
Lemma D.2.
Let $z\in\mathbb{C}$ with $|z|~{}\leq~{}1$. Then $|\mathrm{Im}\left(\arccos(z)\right)|~{}\leq~{}1/n$ whenever $|\arg(z)|~{}\leq~{}1/64n^{2}$.
Proof of Lemma D.2.
Let $re^{i\theta}=z$ and assume $|\theta|~{}\leq~{}1/64n^{2}$. We use the Taylor series for $\arccos(\cdot)$,
$\arccos(re^{i\theta})$ $=\frac{\pi}{2}-\sum_{k=0}^{\infty}a_{k}(re^{i\theta})^{2k+1}$ (For $a_{k}=\frac{(2k)!}{4^{k}(k!)^{2}(2k+1)}$)
$=\frac{\pi}{2}-\sum_{k=0}^{\infty}a_{k}r^{2k+1}e^{i(2k+1)\theta}$ (De Moivre’s Theorem)
$=\frac{\pi}{2}-\sum_{k=0}^{\infty}a_{k}r^{2k+1}\cos((2k+1)\theta)-i\sum_{k=0}^{\infty}a_{k}r^{2k+1}\sin((2k+1)\theta).$ ($e^{i\theta}=\cos\theta+i\sin\theta$)
Therefore,
$\mathrm{Im}\left(\arccos(re^{i\theta})\right)=\sum_{k=0}^{\infty}a_{k}r^{2k+1}\sin((2k+1)\theta).$
Then
$|\mathrm{Im}\left(\arccos(re^{i\theta})\right)|$ $~{}\leq~{}\sum_{k=0}^{\infty}a_{k}|r|^{2k+1}|\sin((2k+1)\theta)|$
$~{}\leq~{}\sum_{k=0}^{\infty}a_{k}|r|^{2k+1}\min(1,(2k+1)|\theta|)$ ($|\sin(x)|~{}\leq~{}\min(|x|,1)$)
$~{}\leq~{}\sum_{k=0}^{\infty}a_{k}\min(1,(2k+1)|\arg(z)|)$ ($|r|~{}\leq~{}1$)
$~{}\leq~{}\sum_{k=0}^{K}a_{k}(2k+1)|\arg(z)|+\sum_{k=K+1}^{\infty}a_{k}.$ (For any arbitrary $K~{}\geq~{}0$)
Now we bound $a_{k}$.
$a_{k}(2k+1)$ $=\frac{(2k)!}{4^{k}(k!)^{2}}$
$~{}\leq~{}\frac{\sqrt{2\pi(2k)}(2k/e)^{2k}\left(1+\frac{1}{2k}\right)}{4^{k}\left(\sqrt{2\pi k}(k/e)^{k}\right)^{2}}$ (Stirling’s Formula )
$=\left(1+\frac{1}{2k}\right)/\sqrt{\pi k}$
$~{}\leq~{}1/\sqrt{k}.$
Therefore we also have that $a_{k}~{}\leq~{}1/k^{3/2}$. Using this (and noting that $a_{0}=1$) we see,
$|\mathrm{Im}\left(\arccos(z)\right)|$ $~{}\leq~{}|\arg(z)|\left(1+\sum_{k=1}^{K}\frac{1}{\sqrt{k}}\right)+\sum_{k=K}^{\infty}\frac{1}{k^{3/2}}$
$~{}\leq~{}4\left(|\arg(z)|\sqrt{K}+\frac{1}{\sqrt{K}}\right)$
$~{}\leq~{}\frac{8}{\sqrt{K}}$ (For $|\arg(z)|~{}\leq~{}1/K$)
$~{}\leq~{}\frac{1}{n}$ (For $K~{}\geq~{}64n^{2}$.)
Therefore, for $|\arg(z)|~{}\leq~{}1/64n^{2}$, we have that
$|\mathrm{Im}(\arccos(z))|~{}\leq~{}1/n.$
${1.5ex}{1.5ex}$ 0.0pt1.5ex1.5ex
Proof of Lemma 4.2.
We bound the coefficients of the Chebyshev polynomial. From Chapter 22 of [3],
$T_{n}(x)=\frac{n}{2}\sum_{m=0}^{\lfloor n/2\rfloor}(-1)^{m}\frac{(n-m-1)!}{m!(n-2m)!}(2x)^{n-2m}.$ (20)
Therefore
$M_{n}(x)=\frac{1}{2^{n-1}}T_{n}(x)=n\sum_{m=0}^{\lfloor n/2\rfloor}(-1)^{m}\frac{(n-m-1)!}{m!(n-2m)!}2^{-2m}x^{n-2m}.$
Let $c_{m}=\frac{(n-m-1)!}{m!(n-2m)!}2^{-2m}$. Then
$\max_{m=0,\dots,n}c_{m}$ $~{}\leq~{}\max_{m=0,\dots,n}{n-m\choose m}4^{-m}$
$~{}\leq~{}\max_{m=0,\dots,n}\left(\frac{(n-m)e}{4m}\right)^{m}$ ((${n\choose k}~{}\leq~{}(ne/k)^{k}$)
$~{}\leq~{}\max_{c\in[0,1]}\left(\frac{(1-c)e}{4c}\right)^{cn}$ (Letting $m=cn$)
$~{}\leq~{}2^{0.3n}.$ ($\max_{c\in[0,1]}((1-c)e/4c)^{c}~{}\leq~{}2^{0.3}$)
${1.5ex}{1.5ex}$ 0.0pt1.5ex1.5ex
Appendix E Plots from Experiments
The details of the experiments are in Section 5, please refer to that for specifics.
E.1 Experiments with Linear Dynamical System Data
Recall we generate the linear data as follows: Sample $(\mathbf{A},\mathbf{B},\mathbf{C})$ with $\mathbf{A}\in{\mathcal{R}}^{300\times 300}$ having eigenvalues $\{z_{j}\}$ drawn uniformly in the complex plane subject to $\mathrm{Im}(z_{j})\leq\tau_{\rm thresh}$ and $L\leq|z_{j}|\leq U$, and $\mathbf{B},\mathbf{C}\in{\mathcal{R}}^{300}$. Then
$\mathbf{x}_{t}=\mathbf{A}\,\mathbf{x}_{t-1}+\mathbf{B}\,\mathbf{u}_{t},\quad\mathbf{y}_{t}=\mathbf{C}\,\mathbf{x}_{t}+\epsilon_{t},\;\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}I).$
We evaluate the following methods: (1) Regression (Alg. 3) , (2) Spectral Filtering (Alg. 4) , (3) DNN Predictor: $n$-layer LSTM with dims $[d_{1},\dots,d_{n}]$, ReLU.
Each method is applied with one of:
  1. 1.
    Baseline: no preconditioning
  2. 2.
    Chebyshev: $\mathbf{c}_{1:n}$ are the coefficients for the $n$th-Chebyshev polynomial
  3. 3.
    Legendre: $\mathbf{c}_{1:n}$ are the coefficients for the $n$th-Legendre polynomial
  4. 4.
    Learned: $\mathbf{c}_{1:n}$ is a parameter learned jointly with the model parameters
We test polynomial degrees $n\in\{2,5,10,20\}$. This choice of degrees shows a rough picture of the impact of $n$. For Chebyshev and Legendre, once the degree is higher than $5-10$ the performance degrades since $\|\mathbf{c}\|_{1}$ gets very large (see our Lemma 4.2 which shows that these coefficients grow exponentially fast).
(a) Regression, $\tau_{\textrm{thresh}}$=0.01
(b) Regression, $\tau_{\textrm{thresh}}$=0.1
(c) Regression, $\tau_{\textrm{thresh}}=0.9$
(d) Spectral Filtering, $\tau_{\textrm{thresh}}=0.01$
(e) Spectral Filtering, $\tau_{\textrm{thresh}}=0.1$
(f) Spectral Filtering, $\tau_{\textrm{thresh}}=0.9$
(g) DNN, $\tau_{\textrm{thresh}}=0.01$
(h) DNN, $\tau_{\textrm{thresh}}=0.1$
(i) DNN, $\tau_{\textrm{thresh}}=0.9$
Figure 2: Absolute prediction error averaged over 200 independent runs with data generated from a linear dynamical system with varying complex threshold.
Next we provide a different plot to show the effect of increased degree of the coefficient. As our theory shows, for Chebyshev polynomials (the same can be shown for Legendre polynomials), the coefficients of the polynomial grow exponentially and therefore the performance gains vanish after the degree is too high. However, learning the optimal coefficients is able to sidestep this issue.
(a) Regression, $\tau_{\textrm{thresh}}$=0.01
(b) Regression, $\tau_{\textrm{thresh}}$=0.1
(c) Regression, $\tau_{\textrm{thresh}}=0.9$
(d) Spectral Filtering, $\tau_{\textrm{thresh}}=0.01$
(e) Spectral Filtering, $\tau_{\textrm{thresh}}=0.1$
(f) Spectral Filtering, $\tau_{\textrm{thresh}}=0.9$
(g) DNN, $\tau_{\textrm{thresh}}=0.01$
(h) DNN, $\tau_{\textrm{thresh}}=0.1$
(i) DNN, $\tau_{\textrm{thresh}}=0.9$
Figure 3: Absolute prediction error averaged over 200 independent runs with data generated from a linear dynamical system with varying complex threshold.
E.2 Nonlinear Data
Recall we generate the nonlinear data as follows: Sample $(\mathbf{A}_{1},\mathbf{B}_{1},\mathbf{C})$ and $(\mathbf{A}_{2},\mathbf{B}_{2})$ with $\mathbf{A}_{i}\in{\mathcal{R}}^{10\times 10}$, $\mathbf{B}_{i},\mathbf{C}\in{\mathcal{R}}^{10}$. Then
$\mathbf{x}_{t}^{(0)}=\mathbf{A}_{1}\mathbf{x}_{t-1}+\mathbf{B}_{1}\mathbf{u}_{t},\;\mathbf{x}_{t}^{(1)}=\sigma\bigl{(}\mathbf{x}_{t}^{(0)}\bigr{)},\;\mathbf{x}_{t}=A_{2}\mathbf{x}_{t}^{(1)}+B_{2}\mathbf{u}_{t},\;\mathbf{y}_{t}=\mathbf{C}\mathbf{x}_{t}+\epsilon_{t}.$
(a) Spectral Filtering, $\tau_{\textrm{thresh}}=0.01$
(b) Spectral Filtering, $\tau_{\textrm{thresh}}=0.1$
(c) Spectral Filtering, $\tau_{\textrm{thresh}}=0.9$
(d) DNN, $\tau_{\textrm{thresh}}=0.01$
(e) DNN, $\tau_{\textrm{thresh}}=0.1$
(f) DNN, $\tau_{\textrm{thresh}}=0.9$
Figure 4: Absolute prediction error averaged over 200 independent runs with data generated from a nonlinear dynamical system with varying complex threshold.
(a) Spectral Filtering, $\tau_{\textrm{thresh}}=0.01$
(b) Spectral Filtering, $\tau_{\textrm{thresh}}=0.1$
(c) Spectral Filtering, $\tau_{\textrm{thresh}}=0.9$
(d) DNN, $\tau_{\textrm{thresh}}=0.01$
(e) DNN, $\tau_{\textrm{thresh}}=0.1$
(f) DNN, $\tau_{\textrm{thresh}}=0.9$
Figure 5: Absolute prediction error averaged over 200 independent runs with data generated from a nonlinear dynamical system with varying complex threshold.
E.3 Data from a DNN
Finally we generate data from a 10-layer sparse neural network which stacks 100-dimensional LSTMs with ReLU nonlinear activations.
(a) 8-layer DNN model
(b) 9-layer DNN model
(c) 10-layer DNN model
Figure 6: Absolute prediction error averaged over 200 independent runs for DNN models of increasing depth.
(a) 8-layer DNN model
(b) 9-layer DNN model
(c) 10-layer DNN model
Figure 7: Absolute prediction error averaged over 200 independent runs for DNN models of increasing depth.